### v1

# Dual-Objective Reinforcement Learning with Novel Hamilton-Jacobi-Bellman Formulations

Abstract
In this work, we extend recent advances that connect Hamilton-Jacobi (HJ) equations with RL to propose two novel value functions for dual-objective satisfaction. Namely, we address: (1) the Reach-Always-Avoid problem – of achieving distinct reward and penalty thresholds – and (2) the Reach-Reach problem – of achieving thresholds of two distinct rewards. In contrast with constrained Markov processes or temporal logic approaches, we are able to derive explicit, tractable Bellman forms in this context by decomposing the problems into reach, avoid, and reach-avoid problems to leverage the recent advances. Moreover, we leverage our analysis to propose a variation of Proximal Policy Optimization, dubbed (DO-HJ-PPO), to solve this class of problems and demonstrate that it bests other baselines in safe-arrival and multi-target tasks, providing a new perspective on constrained decision-making.
I Related Works
Many in learning and autonomy have considered balancing safety and liveness. A few particularly relevant topics are mentioned here. Constrained Markov Decision Processes (CMDPs) are a popular approach to transform constraints into a Lagrangian [1, 2, 3, 4, 5], however, this often requires intricate reward engineering and parameter tuning to balance the combined objective. Similarly, Multi-Objective RL solves the Pareto optimal solution of vector-valued rewards but is not focused on priority-scalarized problems [6, 7, 8]. Goal-Conditioned RL [9, 10, 11] and Linear/Signal Temporal Logic RL [12, 13, 14, 15, 16, 17] generally learn to solve multiple tasks at once, by augmenting the problem to a surrogate problem or automaton, but this is a notoriously challenging approach and there are often no guarantees for the relation between the surrogate and original problem. In this work we are able to derive explicit forms for dual-objective Bellman equations that yield the optimal policy by augmenting the state. This leads to direct approaches for learning the dual-objective values, which is proven to yield improved performance. To do this, we build on traditional dynamic-programming methods that solve Hamilton-Jacobi-Bellman (HJB) equations, and specifically those that connect HJB and RL theories [18, 19, 20, 21].
II Problem Definitions
Consider a Markov decision process (MDP) $\mathcal{M}=\langle\mathcal{{S}},\mathcal{{A}},f\rangle$ consisting of finite state and action spaces $\mathcal{{S}}$ and $\mathcal{{A}}$, and unknown dynamics $f$ that define the deterministic transition $s_{t+1}=f(s_{t},a_{t})$. Let an agent interact with the MDP by selecting an action with policy $\pi:\mathcal{{S}}\to\mathcal{{A}}$ to yield a state trajectory $s_{t}^{\pi}$, i.e. $s_{t+1}^{\pi}=f\left(s_{t}^{\pi},\pi\left(s_{t}^{\pi}\right)\right).$
In this work, we consider the Reach-Always-Avoid (RAA) and Reach-Reach (RR) problems, which both involve balancing two objectives, which are each specified in terms involving a best reward or worst penalty encountered over time. For the RAA problem, let $r,p:\mathcal{{S}}\to\mathbb{R}$ be a reward to be maximized and a penalty to be minimized. For the RR problem, let $r_{1},r_{2}:\mathcal{{S}}\to\mathbb{R}$ be two distinct rewards to be maximized. The agent’s overall objective is to maximize the worst-case outcome between the best-over-time reward and worst-over-time penalty (RAA) or the two best-over-time rewards (RR):
[…]
where $q=-p$. Note that we have introduced this negated penalty $q$ in order to establish a "higher-is-better" convention across subtasks as to compose them in an "and" fashion via the $\min$ operation. For conceptual ease, we recommend the reader still think of trying to minimize the largest-over-time penalty $p$ rather than maximize the smallest-over-time $q$.
Figure 1: DDQN Demonstration of the RAA & RR Problems We compare our novel formulations with previous HJ-RL formulations (RA & R) in a simple grid-world problem with Double-Deep $Q$ Learning. The hazards are highlighted in red, the goal in blue, and trajectories in black (starting at the dot). In both models, the agents actions are limited to left and right or straight and the system flows upwards over time.
As the problem names suggest, these optimization problems are inspired by (but not limited to) tasks involving goal reaching and hazard avoidance. While these problems are thematically distinct, they are mathematically complementary. The respective values corresponding to a policy $\pi$ are
$V_{\textup{RAA}}^{\pi}(s)$ $=\min\left\{\max_{t}r(s_{t}^{\pi}),\>\min_{t}q(s_{t}^{\pi})\right\}$
$V_{\textup{RR}}^{\pi}(s)$ $=\min\left\{\max_{t}r_{1}(s_{t}^{\pi}),\>\max_{t}r_{2}(s_{t}^{\pi})\right\}.$
Observe that the minimums and maximums over time make these values fundamentally different from more typical "sum-of-discounted-reward" objectives considered in RL [22]. Moreover, while both returns involve two objectives, these objectives are combined in worst-case fashion to ensure dual-satisfaction.
III Reachability and Avoidability in RL
Prior works [18, 19] study the reach $V_{\textup{R}}^{\pi}$, avoid $V_{\textup{A}}^{\pi}$, and reach-avoid $V_{\textup{RA}}^{\pi}$ values, respectively defined by
$V_{\textup{R}}^{\pi}(s)$ $=\max_{t}r(s_{t}^{\pi}),$
$V_{\textup{A}}^{\pi}(s)$ $=\min_{t}q(s_{t}^{\pi}),$
$V_{\textup{RA}}^{\pi}(s)$ $=\max_{t}\min\left\{r(s_{t}^{\pi}),\min_{\tau\leq t}q(s_{\tau}^{\pi})\right\},$
for which special Bellman equations [18]. To put these value functions in context, assume the goal $\mathcal{G}$ is the set of states for which $r(s)$ is positive and the hazard $\mathcal{H}$ is the set of states for which $q(s)$ is non-positive. Then $V_{\textup{R}}^{\pi}$, $V_{\textup{A}}^{\pi}$, and $V_{\textup{RA}}^{\pi}$ are positive if and only if $\pi$ causes the agent to eventually reach $\mathcal{G}$, to always avoid $\mathcal{H}$, and to reach $\mathcal{G}$ without hitting $\mathcal{H}$ prior to the reach time, respectively. The Reach-Avoid Bellman Equation (RABE), for example, takes the form [19]
$V_{\textup{RA}}^{*}(s)=\min\left\{\max\left\{\max_{a\in\mathcal{{A}}}V_{\textup{RA}}^{*}\left(f(s,a)\right),r(s)\right\},q(s)\right\},$
and is associated with optimal policy $\pi_{\text{RA}}^{*}(s)$ (without the need for state augmentation, see Section A in the Supplementary Material). This formulation does not naturally induce a contraction, but may be discounted to induce contraction as in [19] by (implicitly) defining $V_{\textup{RA}}^{\gamma}(z)$ for each $\gamma\in[0,1)$ via
$V_{\textup{RA}}^{\gamma}(s)$ $=(1-\gamma)\min\{r(s),q(s)\}\quad+\quad$
$\gamma\min\left\{\max\left\{\max_{a\in\mathcal{{A}}}V_{\textup{RA}}^{\gamma}\left(f(s,a)\right),r(s)\right\},q(s)\right\}.$
These prior value functions and corresponding Bellman equations have proven powerful for these simple reach/avoid/reach-avoid problem formulations. In this work, we generalize these results to the aforementioned broader class of problems.
IV The need for augmenting states
Unlike for the reach/avoid/reach-aviod problems, the goal of choosing a policy $\pi:\mathcal{{S}}\to\mathcal{{A}}$ is inherently flawed for the RR and RAA problems without state augmentation. In considering multiple objectives over an infinite horizon, situations arise in which the optimal action depends on more than the current state, but rather the history the trajectory. An example clarifying the issue is shown in Figure 2.
Figure 2: Examples where a Non-Augmented Policy is Flawed In both MDPs, consider an agent with no memory. (Left) For a deterministic policy based on the current state, the agent can only achieve one target (RR), as the policy must associate the middle state with either of the two possible actions. (Right) In the RAA case, assume the robot must avoid the fire at all costs and would prefer to not encounter the peel, but will do so if needed. The optimal decision for the current state depends on state history, specifically on whether the robot has already reached the target state or not.
To allow the agent to use relevant aspects of its history, we will henceforth consider an augmentation of the MDP with auxiliary variables. A theoretical result in the next section states that this choice of augmentation is sufficient in that no additional information (nor even the use of a stochastic policy) will be able to improve performance under the optimal (deterministic) policy.
IV-A Augmentations
For the RAA problem, we consider an augmentation of the MDP defined by ¯$\overline{\mathcal{M}}=\langle\overline{\mathcal{{S}}},\mathcal{{A}},\bar{f}\rangle$ with augmented states $\overline{\mathcal{{S}}}=\mathcal{{S}}\times\mathcal{{Y}}\times\mathcal{{Z}}$, where $\mathcal{{Y}}:=\{r(s)\mid s\in\mathcal{{S}}\}$ and $\mathcal{{Z}}:=\{q(s)\mid s\in\mathcal{{S}}\}$. Note the actions $\mathcal{{A}}$ remain unchanged. For any initial state $s$, let the augmented states be initialized as $y=r(s)$ and $z=q(s)$, and let the transition $\bar{f}$ of $\overline{\mathcal{M}}$ be given by
$s_{t+1}^{\bar{\pi}}$ $=f\left(s_{t}^{\bar{\pi}},\bar{\pi}\left(s_{t}^{\bar{\pi}},y_{t}^{\bar{\pi}},z_{t}^{\bar{\pi}}\right)\right),$
$y_{t+1}^{\bar{\pi}}$ $=\max\left\{r\left(s_{t+1}^{\bar{\pi}}\right),y_{t}^{\bar{\pi}}\right\},$
$z_{t+1}^{\bar{\pi}}$ $=\min\left\{q\left(s_{t+1}^{\bar{\pi}}\right),z_{t}^{\bar{\pi}}\right\},$
so that $y_{t}$ and $z_{t}$ track the best reward and worst penalty yet seen. Hence, a policy $\bar{\pi}:\overline{\mathcal{{S}}}\to\mathcal{{A}}$ for $\overline{\mathcal{M}}$ may leverage information regarding the history of the trajectory.
For the RR problem, we augment the system similarly, except $\mathcal{{Y}}:=\{r_{1}(s)\mid s\in\mathcal{{S}}\}$, $\mathcal{{Z}}:=\{r_{2}(s)\mid s\in\mathcal{{S}}\}$, and
$s_{t+1}^{\bar{\pi}}$ $=f\left(s_{t}^{\bar{\pi}},\bar{\pi}\left(s_{t}^{\bar{\pi}},y_{t}^{\bar{\pi}},z_{t}^{\bar{\pi}}\right)\right),$
$y_{t+1}^{\bar{\pi}}$ $=\max\left\{r_{1}\left(s_{t+1}^{\bar{\pi}}\right),y_{t}^{\bar{\pi}}\right\},$
$z_{t+1}^{\bar{\pi}}$ $=\max\left\{r_{2}\left(s_{t+1}^{\bar{\pi}}\right),z_{t}^{\bar{\pi}}\right\}.$
V Optimal Policies for RAA and RR by Value Decomposition
We now discuss our first theoretical contributions.
V-A Decomposition of RAA into avoid and reach-avoid problems
Our main theoretical result for the RAA problem shows that we can solve this problem by first solving the avoid problem corresponding to the negated penalty $q(s)$ to obtain the optimal value function $V_{\textup{A}}^{*}(s)$ and then solving a reach-avoid problem with the negated penalty $q(s)$ and a modified reward $r_{\textup{RAA}}(s)$.
Theorem 1.
For all initial states $s\in\mathcal{{S}}$,
$\max_{\bar{\pi}}V_{\textup{RAA}}^{\bar{\pi}}(s)=\max_{\pi}\max_{t}\min\left\{r_{\textup{RAA}}\left(s_{t}^{\pi}\right),\min_{\tau\leq t}q\left(s_{\tau}^{\pi}\right)\right\},$ (1)
where $r_{\textup{RAA}}(s):=\min\left\{r(s),V_{\textup{A}}^{*}(s)\right\}$, with
$V_{\textup{A}}^{*}(s):=\max_{\pi}\min_{t}q\left(s_{t}^{\pi}\right).$
Proof.
See section A of the Supplementary Material. ∎
Corollary 1.
The value function $V_{\textup{RAA}}^{*}(s):=\max_{\bar{\pi}}V_{\textup{RAA}}^{\bar{\pi}}(s)$ satisfies the Bellman equation
$V_{\textup{RAA}}^{*}\left(s\right)=\min\left\{\max\left\{\max_{a\in\mathcal{{A}}}V_{\textup{RAA}}^{*}\left(f(s,a)\right),r_{\textup{RAA}}(s)\right\},q(s)\right\}.$
V-B Decomposition of the RR problem into three reach problems
Our main result for the RR problem shows that we can solve this problem by first solving two reach problems corresponding to the rewards $r_{1}(s)$ and $r_{2}(s)$ to obtain reach value functions $V_{\textup{R1}}^{*}(s)$ and $V_{\textup{R2}}^{*}(s)$, respectively. We then solve a third reach problem with a modified reward $r_{\textup{RR}}(s)$.
Theorem 2.
For all initial states $s\in\mathcal{{S}}$,
$\max_{\bar{\pi}}V_{\textup{RAA}}^{\bar{\pi}}(s)=\max_{\pi}\max_{t}r_{\textup{RR}}\left(s_{t}^{\pi}\right),$ (2)
where
$r_{\textup{RR}}(s):=\max\left\{\min\left\{r_{1}(s),V_{\textup{R2}}^{*}(s)\right\},\min\left\{r_{2}(s),V_{\textup{R1}}^{*}(s)\right\}\right\},$
with
$V_{\textup{R1}}^{*}(s):=\max_{\pi}\max_{t}r_{1}\left(s_{t}^{\pi}\right)\text{ and }V_{\textup{R2}}^{*}(s):=\max_{\pi}\max_{t}r_{2}\left(s_{t}^{\pi}\right).$
Proof.
See section B of the Supplementary Material. ∎
Corollary 2.
The value function $V_{\textup{RR}}^{*}(s):=\max_{\bar{\pi}}V_{\textup{RR}}^{\bar{\pi}}(s)$ satisfies the Bellman equation
$V_{\textup{RR}}^{*}\left(s\right)=\max\left\{\max_{a\in\mathcal{{A}}}V_{\textup{RR}}^{*}\left(f(s,a)\right),r_{\textup{RR}}(s)\right\}.$
V-C Optimality of the augmented problems
The following theoretical result demonstrates that the augmentation is optimal for the original problem in that no further historical information can be used to improve performance.
Theorem 3.
Let $s\in\mathcal{{S}}$. Then
$\max_{\bar{\pi}}V_{\textup{RAA}}^{\bar{\pi}}(s)$ $=\max_{a_{0},a_{1},\dots}\min\left\{\max_{t}r(s_{t}),\min_{t}q(s_{t})\right\},$
and
$\max_{\bar{\pi}}V_{\textup{RR}}^{\bar{\pi}}(s)$ $=\max_{a_{0},a_{1},\dots}\min\left\{\max_{t}r_{1}(s_{t}),\max_{t}r_{2}(s_{t})\right\}$
where $s_{t+1}=f(s_{t},a_{t})$ and $s_{0}=s$.
VI DO-HJ-PPO: Solving RAA and RR with RL
In the previous sections, we demonstrated that the RAA and RR problems can be solved through decomposition of the values into reach, avoid, and reach-avoid problems. Because each of these problems has a corresponding Bellman relation (see [18, 19, 21]), these decompositions allow us to solve the dual-objective problems by minimally modifying existing RL methods. We make a few assumptions in the derivation that would limit performance and generalization, namely, the determinism of the values as well as access to the decomposed values (by solving them beforehand). In this section, we propose relaxations to the RR and RAA theory and devise a custom variant of Proximal Policy Optimization, DO-HJ-PPO, to solve this broader class of problems, and demonstrate its performance.
VI-A Stochastic Reach-Avoid Bellman Equation
It is well known that the most performative RL methods allow for stochastic learning. In [21], the Stochastic Reachability Bellman Equation (SRBE) is described for Reach problems and used to design a specialized PPO algorithm. In this section we proceed by closely following this work, modifying the SRBE into a Stochastic Reach-Avoid Bellman Equation (SRABE). Using Theorems 1 and 2, the SRBE and SRABE offer the necessary tools for designing a PPO variant for solving the RR and RAA problems.
We define $\tilde{V}_{\textup{RAA}}^{\pi}$ to be the solution to the following Bellman equation (SRABE):
$\tilde{V}_{\textup{RAA}}^{\pi}$ $(s)=$
$\mathbb{E}_{a\sim\pi}\left[\min\left\{\max\left\{\tilde{V}_{\textup{RAA}}^{\pi}\left(f(s,a)\right),r_{\textup{RAA}}(s)\right\},q(s)\right\}\right]$
The corresponding action-value function is
$\tilde{Q}_{\textup{RAA}}^{\pi}(s,a)=\min\left\{\max\left\{\tilde{V}_{\textup{RAA}}^{\pi}\left(f(s,a)\right),r_{\textup{RAA}}(s)\right\},q(s)\right\}.$
We define a modification of the dynamics $f$ involving an absorbing state $s_{\infty}$ as follows:
$f^{\prime}(s,a)=\begin{cases}f(s,a)&q\left(f(s,a)\right)<\tilde{V}_{\textup{RAA}}^{\pi}(s)<r_{\textup{RAA}}\left(f(s,a)\right),\\ s_{\infty}&\textup{otherwise}.\end{cases}$
We then have the following proposition:
Proposition 1.
For each $s\in\mathcal{{S}}$ and every $\theta\in\mathbb{R}^{n_{p}}$, we have
$\nabla_{\theta}\tilde{V}_{\textup{RAA}}^{\pi_{\theta}}(s)\propto\mathbb{E}_{s^{\prime}\sim d_{\pi}^{\prime}(s),a\sim\pi_{\theta}}\left[\tilde{Q}_{\textup{RAA}}^{\pi_{\theta}}(s^{\prime},a)\nabla_{\theta}\ln\pi_{\theta}(a|s^{\prime})\right],$
where $d_{\pi}^{\prime}(s)$ is the stationary distribution of the Markov Chain with transition function
$P(s^{\prime}|s)=\sum_{a\in\mathcal{{A}}}\pi(a|s)\left[f^{\prime}(s,\pi(a|s))=s^{\prime}\right],$
with the bracketed term equal to $1$ if the proposition inside is true and 00 otherwise.
Figure 3: Algorithm Comparisons for RR and RAA tasks. We evaluate our method and relevant baselines on 1,000 trajectories for both the Reach-Reach (RR) and Reach-Always-Avoid (RAA) problems for Hopper and F16 environments. In the RR tasks, we compare against a decomposed version of the problem (DSTL) and several variants of CPPO. Our method consistently reaches both target regions with a higher success rate and fewer steps on average. Notably, CPPOv1 and CPPOv2 fail to achieve any successful trajectories in the RR task, whereas CPPOv3 shows improved—but still limited—performance. For the RAA tasks, we compare our approach against Constrained PPO (CPPO) and standard reach-avoid baselines. Our method achieves a higher success rate while requiring a lower average number of steps to reach success.
Following [19], we then define the discounted value and action-value functions with $\gamma\in[0,1)$:
$\tilde{V}_{\textup{RAA}}^{\gamma,\pi}(s)=(1-\gamma)\min\left\{r_{\textup{RAA}}(s),q(s)\right\}$
$\qquad+\gamma\mathbb{E}_{a\sim\pi}\left[\min\left\{\max\left\{\tilde{V}_{\textup{RAA}}^{\gamma,\pi}\left(f(s,a)\right),r_{\textup{RAA}}(s)\right\},q(s)\right\}\right],$
$\tilde{Q}_{\textup{RAA}}^{\gamma,\pi}(s,a)=(1-\gamma)\min\left\{r_{\textup{RAA}}(s),q(s)\right\}$
$\qquad+\gamma\min\left\{\max\left\{\tilde{V}_{\textup{RAA}}^{\gamma,\pi}\left(f(s,a)\right),r_{\textup{RAA}}(s)\right\},q(s)\right\}.$
The PPO advantage function is then given by $\tilde{A}_{\textup{RAA}}^{\pi}=\tilde{Q}_{\textup{RAA}}-\tilde{V}_{\textup{RAA}}$ [23].
VI-B Algorithm
We introduce DO-HJ-PPO in Algorithm 1, a unified PPO-based algorithm for solving the Reach-Always-Avoid (RAA) and Reach-Reach (RR) problems, which builds on the SRABE and SRBE formulations with minimal modifications to the standard PPO framework. Note that in the case of a reset while learning, augmented state variables are set to the initial values described in section IV.
Algorithm 1 : DO-HJ-PPO (Actor-Critic)
1:Define Composed Actor and Critic $\tilde{Q}$
2:Define Decomposed Actor(s) and Critic(s) $\tilde{Q}_{i}$
3:for $k=0,1,\ldots$ do
4:     for $t=0$ to $T-1$ do
5:         Sample trajectories for $\tau_{t}:\{\hat{s}_{t},a_{t},\hat{s}_{t+1}\}$
6:         Define $\tilde{r}(s_{t})$ with Decomposed Critics $\tilde{Q}_{i}(s_{t})$ (Theorems  1 & 2)
7:         Composed Critic update:
$\omega\leftarrow\omega-\beta_{k}\nabla_{\omega}\tilde{Q}(\tau_{t})\cdot\left(\tilde{Q}(\tau_{t})-B^{\gamma}[\tilde{Q};\tilde{r}](\tau_{t})\right)$
8:         Compute Bellman-GAE $A_{HJ}^{\lambda}$ with $B^{\gamma}$
9:         (Standard) update Composed Actor
10:         Decomposed Critic update(s):
$\omega\leftarrow\omega-\beta_{k}\nabla_{\omega}\tilde{Q}_{i}(\tau_{t})\cdot\left(\tilde{Q}_{i}(\tau_{t})-B^{\gamma}_{i}[\tilde{Q}_{i}](\tau_{t})\right)$
11:         Compute Bellman-GAE $A_{i}^{\lambda}$ with $B^{\gamma}_{i}$
12:         (Standard) update Decomposed Actor(s)
13:     end for
14:end for
15:return parameter $\theta$, $\omega$
In Algorithm 1, $B^{\gamma}$ and $B^{\gamma}_{i}$ represent the composed and decomposed Bellman updates for the users choice of problem (RAA or RR). The Bellman update $B^{\gamma}[\tilde{Q};\tilde{r}]$ differs for the RAA task and RR task, depending on the corresponding the reach, avoid, and reach-avoid decomposition. Beyond the modified Bellman updates, the learning proceeds in the nominal PPO fashion, in which the actively-learned decomposed values are used to define the composed value (see Supplementary).
VII Experiments
We first demonstrate the theoretical results (Theorems 1 and 2) through a simple 2D grid-world experiment using Double Deep Q-Networks (DDQN) (Figure 1). Additional experimental details are provided in the Supplementary Material. On the left, we compare the optimal value functions learned under the classic RA formulation with those from the RAA setting. In the RA scenario, trajectories successfully avoid the obstacle but may terminate in regions from which future collisions are inevitable. On the right, we consider a similar environment but with two reward targets. Here, the RR formulation induces trajectories that visit both targets, unlike simple reach tasks in which the agent halts. These qualitative results highlight the behavioral distinctions induced by the RAA and RR objectives compared to their simpler counterparts.
To evaluate the method under more complex and less structured conditions, we extend our analysis to continuous control settings using our algorithm DO-HJ-PPO. Specifically, we apply DO-HJ-PPO to RAA and RR tasks in the Hopper and F16 environments. For the Hopper, two high targets and floor and wall obstacles are defined with respect to its head, and in the F16, the targets are defined by regions to fly through, while obstacles are defined by geofences which create a boxed flight corridor. We compare against both STL (DSTL) and contrained PPO (CPPO) baselines (see supplementary material). Empirically, DO-HJ-PPO performs equivalently at worst and more often at a significantly higher ability, scored in metrics of task success percentage and steps to achieve the task, indicating that DO-HJ-PPO more reliably and rapidly solves the given tasks. These results underscore the challenging nature of composing multiple objectives using traditional baselines while in contrast, our method provides a more robust and direct solution for handling such complex compositional tasks, with less required tuning.
VIII Conclusion
In this brief, we introduced two novel Bellman formulations for new problems (RAA and RR) which generalize those in recent publications. We prove decomposition results for these problems that allow us to break them into simpler Bellman problems, which can then be composed to obtain the value functions and corresponding optimal policies. We use these results to design a PPO-based algorithm for practical solution of RAA and RR. More broadly, this work provides a road-map to extend the range of Bellman formulations that can be solved, via decomposing higher-level problems into lower-level ones. By solving the RAA and RR, we add two new ingredients to the list of solvable problems that can be leveraged toward this end.
Supplementary Material
Contents
-A Proof Notation
Throughout the theoretical sections of this supplement, we use the following notation.
We let =$\mathbb{N}=\{0,1,\dots\}$ be the set of whole numbers.
We let $\mathbb{{A}}$ be the set of maps from $\mathbb{N}$ to $\mathcal{{A}}$. In other words, $\mathbb{{A}}$ is the set of sequences of actions the agent can choose. Given $\mathbf{a}_{1},\mathbf{a}_{2}\in\mathbb{{A}}$, and $\tau\in\mathbb{N}$, we let $[\mathbf{a}_{1},\mathbf{a}_{2}]_{\tau}$ be the element of $\mathbb{{A}}$ for which
$[\mathbf{a}_{1},\mathbf{a}_{2}]_{\tau}(t)=\begin{cases}\mathbf{a}_{1}(t)&t<\tau,\\ \mathbf{a}_{2}(t-\tau)&t\geq\tau.\end{cases}$
Similarly, given $a\in\mathcal{{A}}$ and $\mathbf{a}\in\mathbb{{A}}$, we let $[a,\mathbf{a}]$ be the element of $\mathbb{{A}}$ for which
$[a,\mathbf{a}](t)=\begin{cases}a&t=0,\\ \mathbf{a}(t-1)&t\geq 1.\end{cases}$
Additionally, given $\mathbf{a}\in\mathbb{{A}}$ and $\tau\in\mathbb{N}$, we let $\mathbf{a}|_{\tau}$ be the element of $\mathbb{{A}}$ for which
$\mathbf{a}|_{\tau}(t)=\mathbf{a}(t+\tau)\quad\forall t\in\mathbb{N}.$
The $[\cdot,\cdot]_{\tau}$ operation corresponds to concatenating two action sequences (using only the $0^{\text{th}}$ to $(\tau-1)^{\text{st}}$ elements of the first sequence), the [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] operation corresponds to prepending an action to an action sequence, and the $\cdot|_{\tau}$ operation corresponds to removing the $0^{\text{th}}$ to $(\tau-1)^{\text{st}}$ elements of an action sequence.
We let $\Pi$ be the set of policies $\pi:\mathcal{{S}}\to\mathcal{{A}}$. Given $s\in\mathcal{{S}}$ and $\pi\in\Pi$, we let $\xi_{s}^{\pi}:\mathbb{N}\to\mathcal{{S}}$ be the solution of the evolution equation
$\xi_{s}^{\pi}(t+1)=f\left(\xi_{s}^{\pi}(t),\pi\left(\xi_{s}^{\pi}(t)\right)\right)$
for which $\xi_{s}^{\pi}(0)=s$. In other words, $\xi_{s}^{\pi}(\cdot)$ is the state trajectory over time when the agent begins at state $s$ and follows policy $\pi$.
We will also “overload” this trajectory notation for signals rather than policies: given $\mathbf{a}\in\mathbb{{A}}$, we let $\xi_{s}^{\mathbf{a}}:\mathbb{N}\to\mathcal{{S}}$ be the solution of the evolution equation
$\xi_{s}^{\mathbf{a}}(t+1)=f\left(\xi_{s}^{\mathbf{a}}(t),\mathbf{a}(t)\right)$
for which $\xi_{s}^{\mathbf{a}}(0)=s$. In other words, $\xi_{s}^{\mathbf{a}}(\cdot)$ is the state trajectory over time when the agent begins at state $s$ and follows action sequence $\mathbf{a}$.
-B Proof of RAA Main Theorem
We first define the value functions, $V_{\textup{A}}^{*},\tilde{V}_{\textup{RA}}^{*},V_{\textup{RAA}}^{*}:\mathcal{{S}}\to\mathbb{R}$ by
$V_{\textup{A}}^{*}(s)$ $=\max_{\pi\in\Pi}\min_{\tau\in\mathbb{N}}q\left(\xi_{s}^{\pi}(\tau)\right),$
$\tilde{V}_{\textup{RA}}^{*}(s)$ $=\max_{\pi\in\Pi}\max_{\tau\in\mathbb{N}}\min\left\{r_{\textup{RAA}}\left(\xi_{s}^{\pi}(\tau)\right),\min_{\kappa\leq\tau}q\left(\xi_{s}^{\pi}(\kappa)\right)\right\},$
$V_{\textup{RAA}}^{*}(s)$ $=\max_{\pi\in\Pi}\min\left\{\max_{\tau\in\mathbb{N}}r\left(\xi_{s}^{\pi}(\tau)\right),\min_{\kappa\in\mathbb{N}}q\left(\xi_{s}^{\pi}(\kappa)\right)\right\},$
where $r_{\textup{RAA}}$ is as in Theorem 1.
We next define the value functions, $v_{\textup{A}}^{*},\tilde{v}_{\textup{RA}}^{*},v_{\textup{RAA}}^{*}:\mathcal{{S}}\to\mathbb{R}$, which maximize over action sequences rather than policies:
$v_{\textup{A}}^{*}(s)$ $=\max_{\mathbf{a}\in\mathbb{{A}}}\min_{\tau\in\mathbb{N}}q\left(\xi_{s}^{\mathbf{a}}(\tau)\right),$
$\tilde{v}_{\textup{RA}}^{*}(s)$ $=\max_{\mathbf{a}\in\mathbb{{A}}}\max_{\tau\in\mathbb{N}}\min\left\{r_{\textup{RAA}}\left(\xi_{s}^{\mathbf{a}}(\tau)\right),\min_{\kappa\leq\tau}q\left(\xi_{s}^{\mathbf{a}}(\kappa)\right)\right\},$
$v_{\textup{RAA}}^{*}(s)$ $=\max_{\mathbf{a}\in\mathbb{{A}}}\min\left\{\max_{\tau\in\mathbb{N}}r\left(\xi_{s}^{\mathbf{a}}(\tau)\right),\min_{\kappa\in\mathbb{N}}q\left(\xi_{s}^{\mathbf{a}}(\kappa)\right)\right\},$
Observe that for each $s\in\mathcal{{S}}$,
$v_{\textup{A}}^{*}(s)\geq V_{\textup{A}}^{*}(s),\quad\tilde{v}_{\textup{RA}}^{*}(s)\geq\tilde{V}_{\textup{RA}}^{*}(s),\quad v_{\textup{RAA}}^{*}(s)\geq V_{\textup{RAA}}^{*}(s).$
We now prove a series of lemmas that will be useful in the proof of the main theorem.
Lemma 1.
There is a $\pi\in\Pi$ such that
$v_{\textup{A}}^{*}(s)=\min_{\tau\in\mathbb{N}}q\left(\xi_{s}^{\pi}(\tau)\right)$
for all $s\in\mathcal{{S}}$.
Proof.
Choose $\pi\in\Pi$ such that
$\pi(s)\in\operatorname*{arg\,max}_{a\in\mathcal{{A}}}v_{\textup{A}}^{*}\left(f(s,a)\right)\quad\forall s\in\mathcal{{S}}.$
Fix $s\in\mathcal{{S}}$. Note that for each $\tau\in\mathbb{N}$,
$v_{\textup{A}}^{*}\left(\xi_{s}^{\pi}(\tau+1)\right)$ $=v_{\textup{A}}^{*}\left(f\left(\xi_{s}^{\pi}(\tau),\pi\left(\xi_{s}^{\pi}(\tau)\right)\right)\right)$
$=\max_{a\in\mathcal{{A}}}v_{\textup{A}}^{*}\left(f\left(\xi_{s}^{\pi}(\tau),a\right)\right)$
$=\max_{a\in\mathcal{{A}}}\max_{\mathbf{a}\in\mathbb{{A}}}\min_{\kappa\in\mathbb{N}}q\left(\xi_{f(\xi_{s}^{\pi}(\tau),a)}^{\mathbf{a}}(\kappa)\right)$
$=\max_{a\in\mathcal{{A}}}\max_{\mathbf{a}\in\mathbb{{A}}}\min_{\kappa\in\mathbb{N}}q\left(\xi_{\xi_{s}^{\pi}(\tau)}^{[a,\mathbf{a}]}(\kappa+1)\right)$
$=\max_{\mathbf{a}\in\mathbb{{A}}}\min_{\kappa\in\mathbb{N}}q\left(\xi_{\xi_{s}^{\pi}(\tau)}^{\mathbf{a}}(\kappa+1)\right)$
$\geq\max_{\mathbf{a}\in\mathbb{{A}}}\min_{\kappa\in\mathbb{N}}q\left(\xi_{\xi_{s}^{\pi}(\tau)}^{\mathbf{a}}(\kappa)\right)$
$\geq v_{\textup{A}}^{*}\left(\xi_{s}^{\pi}(\tau)\right).$
It follows by induction that $v_{\textup{A}}^{*}\left(\xi_{s}^{\pi}(\tau)\right)\geq v_{\textup{A}}^{*}\left(\xi_{s}^{\pi}(0)\right)$ for all $\tau\in\mathbb{N}$, so that
$v_{\textup{A}}^{*}(s)\geq\min_{\tau\in\mathbb{N}}q\left(\xi_{s}^{\pi}(\tau)\right)\geq\min_{\tau\in\mathbb{N}}v_{\textup{A}}^{*}\left(\xi_{s}^{\pi}(\tau)\right)=v_{\textup{A}}^{*}\left(\xi_{s}^{\pi}(0)\right)=v_{\textup{A}}^{*}(s).$
Corollary 3.
For all $s\in\mathcal{{S}}$, we have $V_{\textup{A}}^{*}(s)=v_{\textup{A}}^{*}(s)$.
Lemma 2.
There is a $\pi\in\Pi$ such that
$\tilde{v}_{\textup{RA}}^{*}(s)=\max_{\tau\in\mathbb{N}}\min\left\{r_{\textup{RAA}}\left(\xi_{s}^{\pi}(\tau)\right),\min_{\kappa\leq\tau}q\left(\xi_{s}^{\pi}(\kappa)\right)\right\}$
for all $s\in\mathcal{{S}}$.
Proof.
First, let us note that in this proof we will use the standard conventions that
$\max\varnothing=-\infty\quad\text{and}\quad\min\varnothing=+\infty.$
We next introduce some notation. First, for convenience, we set $v^{*}=\tilde{v}_{\textup{RA}}^{*}$ and $V^{*}=\tilde{V}_{\textup{RA}}^{*}$. Given $s\in\mathcal{{S}}$ and $\mathbf{a}\in\mathbb{{A}}$, we write
$v^{\mathbf{a}}(s)=\max_{\tau\in\mathbb{N}}\min\left\{r_{\textup{RAA}}\left(\xi_{s}^{\mathbf{a}}(\tau)\right),\min_{\kappa\leq\tau}q\left(\xi_{s}^{\mathbf{a}}(\kappa)\right)\right\}.$
Similarly, given $s\in\mathcal{{S}}$ and $\pi\in\Pi$, we write
$V^{\pi}(s)=\max_{\tau\in\mathbb{N}}\min\left\{r_{\textup{RAA}}\left(\xi_{s}^{\pi}(\tau)\right),\min_{\kappa\leq\tau}q\left(\xi_{s}^{\pi}(\kappa)\right)\right\}.$
Then
$V^{*}(s)=\max_{\pi\in\Pi}\max_{\tau\in\mathbb{N}}\min\left\{r_{\textup{RAA}}\left(\xi_{s}^{\pi}(\tau)\right),\min_{\kappa\leq\tau}q\left(\xi_{s}^{\pi}(\kappa)\right)\right\}=\max_{\pi\in\Pi}V^{\pi}(s),$
and
$v^{*}(s)=\max_{\mathbf{a}\in\mathbb{{A}}}\max_{\tau\in\mathbb{N}}\min\left\{r_{\textup{RAA}}\left(\xi_{s}^{\mathbf{a}}(\tau)\right),\min_{\kappa\leq\tau}q\left(\xi_{s}^{\mathbf{a}}(\kappa)\right)\right\}=\max_{\mathbf{a}\in\mathbb{{A}}}v^{\mathbf{a}}(s).$
It is immediate that $v^{*}(s)\geq V^{*}(s)$ for each $s\in\mathcal{{S}}$, so it suffices to show the reverse inequality. Toward this end, it suffices to show that there is a $\pi\in\Pi$ for which $V^{\pi}(s)=v^{*}(s)$ for each $s\in\mathcal{{S}}$. Indeed, in this case, $V^{*}(s)\geq V^{\pi}(s)=v^{*}(s)$.
We now construct the desired policy $\pi$. Let $\alpha_{0}=+\infty$, $S_{0}=\varnothing$, and $v_{0}^{*}:\mathcal{{S}}\to\mathbb{R}\cup\{-\infty\},s\mapsto-\infty$. We recursively define $\alpha_{t}\in\mathbb{R}$, $S_{t}\subseteq\mathcal{{S}}$, and $v_{t}^{*}:\mathcal{{S}}\to\mathbb{R}\cup\{-\infty\}$ for $t=1,2,\dots$ by
$\alpha_{t+1}=\max_{s\in\mathcal{{S}}\setminus S_{t}}\min\left\{\max\left\{r_{\textup{RAA}}(s),\max_{a\in\mathcal{{A}}}v_{t}^{*}\left(f(s,a)\right)\right\},q(s)\right\},$ (3)
$S_{t+1}=S_{t}\cup\left\{s\in\mathcal{{S}}\setminus S_{t}~{}\biggl{\lvert}~{}\min\left\{\max\left\{r_{\textup{RAA}}(s),\max_{a\in\mathcal{{A}}}v_{t}^{*}\left(f(s,a)\right)\right\},q(s)\right\}=\alpha_{t+1}\right\},$ (4)
$v_{t+1}^{*}(s)=\begin{cases}v_{t}^{*}(s)&s\in S_{t},\\ \alpha_{t+1}&s\in S_{t+1}\setminus S_{t},\\ -\infty&s\in\mathcal{{S}}\setminus S_{t+1}.\end{cases}$ (5)
From (4) it follows that
$S_{0}\subseteq S_{1}\subseteq S_{2}\subseteq\dots,$ (6)
which together with (3) shows that
$\alpha_{0}\geq\alpha_{1}\geq\alpha_{2}\geq\dots.$ (7)
Also, whenever $\mathcal{{S}}\setminus S_{t}$ is non-empty, the set being appended to $S_{t}$ in (4) is non-empty so
$\bigcup_{t=0}^{\infty}S_{t}=\mathcal{{S}}.$ (8)
For each $s\in\mathcal{{S}}$, let $\sigma(s)$ be the smallest $t\in\mathbb{N}$ for which $s\in S_{t}$. We choose the policy $\pi\in\Pi$ of interest by insisting
$\pi(s)\in\operatorname*{arg\,max}_{a\in\mathcal{{A}}}v_{\sigma(s)-1}^{*}(f(s,a))\quad\forall s\in\mathcal{{S}}.$ (9)
In the remainder of the proof, we show that $V^{\pi}(s)=v^{*}(s)$ for each $s\in\mathcal{{S}}$ by induction. Let $n\in\mathbb{N}$ and suppose the following induction assumptions hold:
$V^{\pi}(s)$ $=v^{*}(s)=v_{n}^{*}(s)\geq\alpha_{n}\quad\forall s\in S_{n},$ (10)
$v^{*}(s^{\prime})$ $\leq\alpha_{n}\quad\forall s^{\prime}\in\mathcal{{S}}\setminus S_{n}.$ (11)
Note that the above hold trivially when $n=0$ since $S_{0}=\varnothing$ and $\alpha_{0}=+\infty$. Fix some particular $y\in S_{n+1}$ and some $z\in\mathcal{{S}}\setminus S_{n+1}$. We must show that
$V^{\pi}(y)$ $=v^{*}(y)=v_{n+1}^{*}(y)\geq\alpha_{n+1},$ (12)
$v^{*}(z)$ $\leq\alpha_{n+1}.$ (13)
In this case, induction then shows that $V^{\pi}(s)=v^{*}(s)$ for all $s\in\cup_{n=0}^{\infty}S_{t}$. Since this union is equal to $\mathcal{{S}}$ by (8), the desired result then follows.
To show (12)-(13), we first demonstrate the following three claims.
  1. 1.
    Let $x\in\mathcal{{S}}$ and $w\in\mathcal{{A}}$ be such that $f(x,w)\in S_{n}$ and $q(x)\geq\alpha_{n+1}$. We claim $x\in S_{n+1}$.
    We can assume $x\notin S_{n}$, for otherwise the claim follows immediately from (6). Since $f(x,w)\in S_{n}$, we have $v_{n}^{*}\left(f(x,w)\right)\geq\alpha_{n}$ by (10). Thus
    $\alpha_{n+1}$ $\geq\min\left\{\max\left\{r_{\textup{RAA}}(x),\max_{a\in\mathcal{{A}}}v_{n}^{*}\left(f(x,a)\right)\right\},q(x)\right\}$
    $\geq\min\left\{\max\{r_{\textup{RAA}}(x),\alpha_{n}\},\alpha_{n+1}\right\}$
    $=\alpha_{n+1},$
    where the first inequality follows from (3), and the equality follows from (7). Thus
    $\alpha_{n+1}=\min\left\{\max\left\{r_{\textup{RAA}}(x),\max_{a\in\mathcal{{A}}}v_{n}^{*}\left(f(x,a)\right)\right\},q(x)\right\},$
    so the claim follows from (4).
  2. 2.
    Let $x\in S_{n+1}\setminus S_{n}$ and $w\in\mathcal{{A}}$ be such that $f(x,w)\in S_{n}$. We claim that
    $V^{\pi}(x)=v^{*}(x)=\alpha_{n+1}.$ (14)
    To show this claim, we will make use of the dynamic programming principle
    $v^{\mathbf{a}}(s)=\min\left\{\max\left\{r_{\textup{RAA}}(s),v^{\mathbf{a}|_{1}}\left(f(s,\mathbf{a}(0))\right)\right\},q(s)\right\},\quad\forall s\in\mathcal{{S}},\mathbf{a}\in\mathbb{{A}},$
    from which it follows that
    $V^{\pi}(s)=\min\left\{\max\left\{r_{\textup{RAA}}(s),V^{\pi}\left(f(s,\pi(s))\right)\right\},q(s)\right\},\quad\forall s\in\mathcal{{S}},$ (15)
    and
    $v^{*}(s)=\min\left\{\max\left\{r_{\textup{RAA}}(s),\max_{a\in\mathcal{{A}}}v^{*}\left(f(s,a)\right)\right\},q(s)\right\},\quad\forall s\in\mathcal{{S}}.$ (16)
    Since $x\in S_{n+1}\setminus S_{n}$, then $\sigma(x)=n+1$ by definition of $\sigma$, so $\pi(x)\in\operatorname*{arg\,max}_{a\in\mathcal{{A}}}v_{n}^{*}(f(x,a))$ by (9). Thus
    $v_{n}^{*}\left(f(x,\pi(x))\right)=\max_{a\in\mathcal{{A}}}v_{n}^{*}\left(f(x,a)\right).$ (17)
    But then
    $v_{n}^{*}\left(f(x,\pi(x))\right)\geq v_{n}^{*}\left(f(x,w)\right)\geq\alpha_{n}\geq\alpha_{n+1}>-\infty,$
    where the second inequality comes from (10), the third comes from (7), and the final inequality comes from (3) ($\mathcal{{S}}\setminus S_{n}$ is non-empty because $x\in\mathcal{{S}}\setminus S_{n}$). Thus $f(x,\pi(x))\in S_{n}$ by (5). It then follows from (10) that
    $V^{\pi}\left(f(x,\pi(x))\right)=v^{*}\left(f(x,\pi(x))\right)=v_{n}^{*}\left(f(x,\pi(x))\right).$ (18)
    Now, observe that for all $s\in S_{n}$ and $s^{\prime}\in\mathcal{{S}}\setminus S_{n}$,
    $v^{*}(s)=v_{n}^{*}(s)\geq\alpha_{n}\geq v^{*}(s^{\prime})\geq-\infty=v_{n}^{*}(s^{\prime}),$ (19)
    where the first equality and inequality are from (10), the second inequality is from (11), and the final equality is from (5). Moreover, $f(x,a)\in S_{n}$ for at least one $a$ (in particular $a=w$). Letting $\mathcal{{A}}^{\prime}=\left\{a\in\mathcal{{A}}\mid f(x,a)\in S_{n}\right\}$, it follows from (19) that
    $\max_{a\in\mathcal{{A}}}v^{*}\left(f(x,a)\right)=\max_{a\in\mathcal{{A}}^{\prime}}v^{*}\left(f(x,a)\right)=\max_{a\in\mathcal{{A}}^{\prime}}v_{n}^{*}\left(f(x,a)\right)=\max_{a\in\mathcal{{A}}}v_{n}^{*}\left(f(x,a)\right).$ (20)
    From (17)-(20) we have
    $V^{\pi}\left(f(x,\pi(x))\right)=\max_{a\in\mathcal{{A}}}v^{*}\left(f(x,a)\right)=\max_{a\in\mathcal{{A}}}v_{n}^{*}\left(f(x,a)\right).$ (21)
    Now observe that
    $V^{\pi}(x)$ $=\min\left\{\max\left\{r_{\textup{RAA}}(x),V^{\pi}\left(f(x,\pi(x))\right)\right\},q(x)\right\},$
    $v^{*}(x)$ $=\min\left\{\max\left\{r_{\textup{RAA}}(x),\max_{a\in\mathcal{{A}}}v^{*}\left(f(x,a)\right)\right\},q(x)\right\},$
    $\alpha_{n+1}$ $=\min\left\{\max\left\{r_{\textup{RAA}}(x),\max_{a\in\mathcal{{A}}}v_{n}^{*}\left(f(x,a)\right)\right\},q(x)\right\},$
    where the first equation is from (15), the second is from (16), and the third is from (4). But then (14) follows from the above equations together with (21).
  3. 3.
    Let $x\in\mathcal{{S}}\setminus S_{n}$. We claim that $v^{*}(x)\leq\alpha_{n+1}$. Suppose otherwise. Then we can choose $\mathbf{a}\in\mathbb{{A}}$ and $\tau\in\mathbb{N}$ such that
    $\min\left\{r_{\textup{RAA}}\left(\xi_{x}^{\mathbf{a}}(\tau)\right),\min_{\kappa\leq\tau}q\left(\xi_{x}^{\mathbf{a}}(\kappa)\right)\right\}>\alpha_{n+1}.$ (22)
    It follows that $\xi_{x}^{\mathbf{a}}(\tau)\in S_{n}$, for otherwise
    $\alpha_{n+1}\geq\min\left\{r_{\textup{RAA}}(\xi_{x}^{\mathbf{a}}(\tau)),q(\xi_{x}^{\mathbf{a}}(\tau))\right\}$
    by (3), creating a contradiction.
    So $x\notin S_{n}$ and $\xi_{x}^{\mathbf{a}}(\tau)\in S_{n}$, indicating that there is some $\theta\in\{0,\dots,\tau-1\}$ such that $\xi_{x}^{\mathbf{a}}(\theta)\notin S_{n}$ and $f\left(\xi_{x}^{\mathbf{a}}(\theta),\mathbf{a}(\theta)\right)=\xi_{x}^{\mathbf{a}}(\theta+1)\in S_{n}$. Moreover, $q\left(\xi_{x}^{\mathbf{a}}(\theta)\right)>\alpha_{n+1}$ by (22). It follows from claim 1 that $\xi_{x}^{\mathbf{a}}(\theta)\in S_{n+1}$.
    But then it follows from claim 2 that $v^{*}\left(\xi_{x}^{\mathbf{a}}(\theta)\right)=\alpha_{n+1}.$ However,
    $v^{*}\left(\xi_{x}^{\mathbf{a}}(\theta)\right)$ $\geq\min\left\{r_{\textup{RAA}}\left(\xi_{\xi_{x}^{\mathbf{a}}(\theta)}^{\mathbf{a}}(\tau-\theta)\right),\min_{\kappa\leq\tau-\theta}q\left(\xi_{\xi_{x}^{\mathbf{a}}(\theta)}^{\mathbf{a}}(\kappa)\right)\right\}$
    $=\min\left\{r_{\textup{RAA}}\left(\xi_{x}^{\mathbf{a}}(\tau-\theta+\theta)\right),\min_{\kappa\leq\tau-\theta}q\left(\xi_{x}^{\mathbf{a}}(\kappa+\theta)\right)\right\}$
    $=\min\left\{r_{\textup{RAA}}\left(\xi_{x}^{\mathbf{a}}(\tau)\right),\min_{\kappa\in\{\theta,\theta+1,\dots,\tau\}}q\left(\xi_{x}^{\mathbf{a}}(\kappa)\right)\right\}$
    $>\alpha_{n+1},$
    giving the desired contradiction.
Having established these claims, we return to proving (12) and (13) hold. In fact, (13) follows immediately from claim 3, so we actually only need to show (12).
[…]
$v^{*}(y)\geq V^{\pi}(y)\geq\min\left\{r_{\textup{RAA}}(y),q(y)\right\}=\alpha_{n+1}\geq v^{*}(y),$
where the final inequality follows from claim 3. This completes the proof. ∎
Corollary 4.
For all $s\in\mathcal{{S}}$, we have $\tilde{V}_{\textup{RA}}^{*}(s)=\tilde{v}_{\textup{RA}}^{*}(s)$.
Lemma 3.
Let $F:\mathbb{{A}}\times\mathbb{N}\to\mathbb{R}$. Then
$\sup_{\mathbf{a}\in\mathbb{{A}}}\sup_{\tau\in\mathbb{N}}\sup_{\mathbf{a}^{\prime}\in\mathbb{{A}}^{\prime}}F\left([\mathbf{a},\mathbf{a}^{\prime}]_{\tau},\tau\right)=\sup_{\mathbf{a}\in\mathbb{{A}}}\sup_{\tau\in\mathbb{N}}F\left(\mathbf{a},\tau\right).$ (23)
Proof.
We proceed by showing both inequalities corresponding to (23) hold.
  • ($\geq$)
    Given any $\mathbf{a}\in\mathbb{{A}}$ and $\tau\in\mathbb{N}$, we have $\sup_{\mathbf{a}^{\prime}\in\mathbb{{A}}^{\prime}}F\left([\mathbf{a},\mathbf{a}^{\prime}]_{\tau},\tau\right)\geq F\left(\mathbf{a},\tau\right)$. Taking the suprema over $\mathbf{a}\in\mathbb{{A}}$ and $\tau\in\mathbb{N}$ on both sides of this inequality gives the desired result.
  • ($\leq$)
    Given any $\mathbf{a}\in\mathbb{{A}}$ and $\tau\in\mathbb{N}$, we have
    $\sup_{\mathbf{a}^{\prime}\in\mathbb{{A}}^{\prime}}F\left([\mathbf{a},\mathbf{a}^{\prime}]_{\tau},\tau\right)\leq\sup_{\mathbf{a}^{\prime\prime}\in\mathbb{{A}}}F\left(\mathbf{a}^{\prime\prime},\tau\right),$
    so that the result follows from taking the suprema over $\mathbf{a}\in\mathbb{{A}}$ and $\tau\in\mathbb{N}$ on both sides of this inequality.
Lemma 4.
For each $s\in\mathcal{{S}}$,
$v_{\textup{RAA}}^{*}(s)=\tilde{v}_{\textup{RA}}^{*}(s).$
Proof.
For each $s\in\mathcal{{S}}$, we have
$\tilde{v}_{\textup{RA}}^{*}(s)$ $=\max_{\mathbf{a}\in\mathbb{{A}}}\max_{\tau\in\mathbb{N}}\min\left\{r_{\textup{RAA}}\left(\xi_{s}^{\mathbf{a}}(\tau)\right),\min_{\kappa\leq\tau}q\left(\xi_{s}^{\mathbf{a}}(\kappa)\right)\right\}$ (24)
$=\max_{\mathbf{a}\in\mathbb{{A}}}\max_{\tau\in\mathbb{N}}\min\left\{r\left(\xi_{s}^{\mathbf{a}}(\tau)\right),v_{\textup{A}}^{*}\left(\xi_{s}^{\mathbf{a}}(\tau)\right),\min_{\kappa\leq\tau}q\left(\xi_{s}^{\mathbf{a}}(\kappa)\right)\right\}$ (25)
$=\max_{\mathbf{a}\in\mathbb{{A}}}\max_{\tau\in\mathbb{N}}\min\left\{r\left(\xi_{s}^{\mathbf{a}}(\tau)\right),\max_{\mathbf{a}^{\prime}\in\mathbb{{A}}}\min_{\kappa^{\prime}\in\mathbb{N}}q\left(\xi_{\xi_{s}^{\mathbf{a}}(\tau)}^{\mathbf{a}^{\prime}}(\kappa^{\prime})\right),\min_{\kappa\leq\tau}q\left(\xi_{s}^{\mathbf{a}}(\kappa)\right)\right\}$
$=\max_{\mathbf{a}\in\mathbb{{A}}}\max_{\tau\in\mathbb{N}}\min\left\{r\left(\xi_{s}^{\mathbf{a}}(\tau)\right),\max_{\mathbf{a}^{\prime}\in\mathbb{{A}}}\min_{\kappa^{\prime}\in\mathbb{N}}q\left(\xi_{s}^{[\mathbf{a},\mathbf{a}^{\prime}]_{\tau}}(\tau+\kappa^{\prime})\right),\min_{\kappa\leq\tau}q\left(\xi_{s}^{\mathbf{a}}(\kappa)\right)\right\}$
$=\max_{\mathbf{a}\in\mathbb{{A}}}\max_{\tau\in\mathbb{N}}\max_{\mathbf{a}^{\prime}\in\mathbb{{A}}}\min\left\{r\left(\xi_{s}^{\mathbf{a}}(\tau)\right),\min_{\kappa^{\prime}\in\mathbb{N}}q\left(\xi_{s}^{[\mathbf{a},\mathbf{a}^{\prime}]_{\tau}}(\tau+\kappa^{\prime})\right),\min_{\kappa\leq\tau}q\left(\xi_{s}^{\mathbf{a}}(\kappa)\right)\right\}$
$=\max_{\mathbf{a}\in\mathbb{{A}}}\max_{\tau\in\mathbb{N}}\max_{\mathbf{a}^{\prime}\in\mathbb{{A}}}\min\left\{r\left(\xi_{s}^{[\mathbf{a},\mathbf{a}^{\prime}]_{\tau}}(\tau)\right),\min_{\kappa^{\prime}\in\mathbb{N}}q\left(\xi_{s}^{[\mathbf{a},\mathbf{a}^{\prime}]_{\tau}}(\tau+\kappa^{\prime})\right),\min_{\kappa\leq\tau}q\left(\xi_{s}^{[\mathbf{a},\mathbf{a}^{\prime}]_{\tau}}(\kappa)\right)\right\}$ (26)
$=\max_{\mathbf{a}\in\mathbb{{A}}}\max_{\tau\in\mathbb{N}}\min\left\{r\left(\xi_{s}^{\mathbf{a}}(\tau)\right),\min_{\kappa^{\prime}\in\mathbb{N}}q\left(\xi_{s}^{\mathbf{a}}(\tau+\kappa^{\prime})\right),\min_{\kappa\leq\tau}q\left(\xi_{s}^{\mathbf{a}}(\kappa)\right)\right\}$ (27)
$=\max_{\mathbf{a}\in\mathbb{{A}}}\max_{\tau\in\mathbb{N}}\min\left\{r\left(\xi_{s}^{\mathbf{a}}(\tau)\right),\min_{\kappa\in\mathbb{N}}q\left(\xi_{s}^{\mathbf{a}}(\kappa)\right)\right\}$
$=\max_{\mathbf{a}\in\mathbb{{A}}}\min\left\{\max_{\tau\in\mathbb{N}}r\left(\xi_{s}^{\mathbf{a}}(\tau)\right),\min_{\kappa\in\mathbb{N}}q\left(\xi_{s}^{\mathbf{a}}(\kappa)\right)\right\}$
$=v_{\textup{RAA}}^{*}(s),$
where the equality between (24) and (25) follows from Corollary 3, and where the equality between (26) and (27) follows from Lemma 3. ∎
Before the next lemma, we need to introduce two last pieces of notation. First, we let $\overline{\Pi}$ be the set of augmented policies $\bar{\pi}:\mathcal{{S}}\times\mathcal{{Y}}\times\mathcal{{Z}}\to\mathcal{{A}}$, where
$\mathcal{{Y}}=\left\{r(s)\mid s\in\mathcal{{S}}\right\}\quad\text{and}\quad\mathcal{{Z}}=\left\{q(s)\mid s\in\mathcal{{S}}\right\}.$
Next, given $s\in\mathcal{{S}}$, $y\in\mathcal{{Y}}$, $z\in\mathcal{{Z}}$, and $\bar{\pi}\in\overline{\Pi}$, we let $\bar{\xi}_{s}^{\bar{\pi}}:\mathbb{N}\to\mathcal{{S}}$, $\bar{\eta}_{s}^{\bar{\pi}}:\mathbb{N}\to\mathcal{{Y}}$, and $\bar{\zeta}_{s}^{\bar{\pi}}:\mathbb{N}\to\mathcal{{Z}}$, be the solution of the evolution
$\bar{\xi}_{s}^{\bar{\pi}}(t+1)$ $=f\left(\bar{\xi}_{s}^{\bar{\pi}}(t),\bar{\pi}\left(\bar{\xi}_{s}^{\bar{\pi}}(t),\bar{\eta}_{s}^{\bar{\pi}}(t),\bar{\zeta}_{s}^{\bar{\pi}}(t)\right)\right),$
$\bar{\eta}_{s}^{\bar{\pi}}(t+1)$ $=\max\left\{r\left(\bar{\xi}_{s}^{\bar{\pi}}(t+1)\right),\bar{\eta}_{s}^{\bar{\pi}}(t)\right\},$
$\bar{\zeta}_{s}^{\bar{\pi}}(t+1)$ $=\min\left\{q\left(\bar{\xi}_{s}^{\bar{\pi}}(t+1)\right),\bar{\zeta}_{s}^{\bar{\pi}}(t)\right\},$
for which $\bar{\xi}_{s}^{\bar{\pi}}(0)=s$, $\bar{\eta}_{s}^{\bar{\pi}}(0)=r(s)$, and $\bar{\zeta}_{s}^{\bar{\pi}}(0)=q(s)$.
Lemma 5.
There is a $\bar{\pi}\in\overline{\Pi}$ such that
$v_{\textup{RAA}}^{*}(s)=\min\left\{\max_{\tau\in\mathbb{N}}r\left(\bar{\xi}_{s}^{\bar{\pi}}(\tau)\right),\min_{\tau\in\mathbb{N}}q\left(\bar{\xi}_{s}^{\bar{\pi}}(\tau)\right)\right\}$ (28)
for all $s\in\mathcal{{S}}$.
Proof.
By Lemmas 1 and 2 together with Corollary 3, we can choose $\pi,\theta\in\Pi$ such that
$\tilde{v}_{\textup{RA}}^{*}(s)$ $=\max_{\tau\in\mathbb{N}}\min\left\{r\left(\xi_{s}^{\pi}(\tau)\right),v_{\textup{A}}^{*}\left(\xi_{s}^{\pi}(\tau)\right),\min_{\kappa\leq\tau}q\left(\xi_{s}^{\pi}(\kappa)\right)\right\}\quad\forall s\in\mathcal{{S}},$
$v_{\textup{A}}^{*}(s)$ $=\min_{\tau\in\mathbb{N}}q\left(\xi_{s}^{\theta}(\tau)\right)\quad\forall s\in\mathcal{{S}}.$
We introduce some useful notation we will use throughout the rest of the proof. For each $s\in\mathcal{{S}}$, let $[s]^{+}=f(s,\pi(s))$, $[y]^{+}_{s}=\max\{y,r\left([s]^{+}\right)\}$, $[z]^{+}_{s}=\min\{z,q\left([s]^{+}\right)\}$.
We define an augmented policy $\bar{\pi}\in\overline{\Pi}$ by
$\bar{\pi}(s,y,z)=\begin{cases}\pi(s)&\min\{[y]_{s}^{+},[z]_{s}^{+},v_{\textup{A}}^{*}([s]^{+})\}\geq\min\{y,z,v_{\textup{A}}^{*}(s)\},\\ \theta(s)&\text{otherwise}.\end{cases}$
Now fix some $s\in\mathcal{{S}}$. For all $t\in\mathbb{N}$, set $\bar{x}_{t}=\bar{\xi}_{s}^{\bar{\pi}}(t)$, $\bar{y}_{t}=\bar{\eta}_{s}^{\bar{\pi}}(t)=\max_{\tau\leq t}r(\bar{x}_{\tau})$, and $\bar{z}_{t}=\bar{\zeta}_{s}^{\bar{\pi}}(t)=\min_{\tau\leq t}q(\bar{x}_{\tau})$, and also set $x^{\circ}_{t}=\xi_{s}^{\pi}(t)$, $y^{\circ}_{t}=\max_{\tau\leq t}r(x^{\circ}_{\tau})$, and $z^{\circ}_{t}=\min_{\tau\leq t}q(x^{\circ}_{\tau})$.
First, assume that $t$ is such that $\min\{[\bar{y}_{t}]_{\bar{x}_{t}}^{+},[\bar{z}_{t}]_{\bar{x}_{t}}^{+},v_{\textup{A}}^{*}([\bar{x}_{t}]^{+})\}<\min\{\bar{y}_{t},\bar{z}_{t},v_{\textup{A}}^{*}(\bar{x}_{t})\}$. In this case, $\bar{\pi}(\bar{x}_{t},\bar{y}_{t},\bar{z}_{t})=\theta(\bar{x}_{t})$, so that
$\min\{\bar{z}_{t},v_{\textup{A}}^{*}(\bar{x}_{t})\}=\min\{\bar{z}_{t+1},v_{\textup{A}}^{*}(\bar{x}_{t+1})\}$
by our choice of $\theta$. Since $\bar{y}_{t}$ is non-decreasing in $t$, thus have
$\min\{\bar{y}_{t},\bar{z}_{t},v_{\textup{A}}^{*}(\bar{x}_{t})\}\leq\min\{\bar{y}_{t+1},\bar{z}_{t+1},v_{\textup{A}}^{*}(\bar{x}_{t+1})\}.$
Next, assume that $t$ is such that $\min\{[\bar{y}_{t}]_{\bar{x}_{t}}^{+},[\bar{z}_{t}]_{\bar{x}_{t}}^{+},v_{\textup{A}}^{*}([\bar{x}_{t}]^{+})\}\geq\min\{\bar{y}_{t},\bar{z}_{t},v_{\textup{A}}^{*}(\bar{x}_{t})\}$. In this case, we have that $\bar{\pi}(\bar{x}_{t},\bar{y}_{t},\bar{z}_{t})=\pi(\bar{x}_{t})$, so
$\min\{\bar{y}_{t},\bar{z}_{t},v_{\textup{A}}^{*}(\bar{x}_{t})\}\leq\min\{[\bar{y}_{t}]_{\bar{x}_{t}}^{+},[\bar{z}_{t}]_{\bar{x}_{t}}^{+},v_{\textup{A}}^{*}([\bar{x}_{t}]^{+})\}=\min\{\bar{y}_{t+1},\bar{z}_{t+1},v_{\textup{A}}^{*}(\bar{x}_{t+1})\}.$
It thus follows from these two cases that $\min\{\bar{y}_{t},\bar{z}_{t},v_{\textup{A}}^{*}(\bar{x}_{t})\}$ is non-decreasing in $t$. Let
$T=\min\left\{t\in\mathbb{N}\mid\min\{[\bar{y}_{t}]_{\bar{x}_{t}}^{+},[\bar{z}_{t}]_{\bar{x}_{t}}^{+},v_{\textup{A}}^{*}([\bar{x}_{t}]^{+})\}<\min\{\bar{y}_{t},\bar{z}_{t},v_{\textup{A}}^{*}(\bar{x}_{t})\}\right\}.$
There are again two cases:
  • ($T<\infty$)
    In this case, $\bar{\pi}(\bar{x}_{t},\bar{y}_{t},\bar{z}_{t})=\pi(\bar{x}_{t})$ for $t<T$. Then $\bar{x}_{t}=x^{\circ}_{t}$, $\bar{y}_{t}=y^{\circ}_{t}$, and $\bar{z}_{t}=z^{\circ}_{t}$ for all $t\leq T$. It follows that $[\bar{x}_{t}]^{+}=x^{\circ}_{t+1}$, $[\bar{y}_{t}]_{\bar{x}_{t}}^{+}=y^{\circ}_{t+1}$, and $[\bar{z}_{t}]_{\bar{x}_{t}}^{+}=z^{\circ}_{t+1}$ for all $t\leq T$. Thus by definition of $T$,
    $\min\left\{y^{\circ}_{t+1},z^{\circ}_{t+1},v_{\textup{A}}^{*}\left(x^{\circ}_{t+1}\right)\right\}\geq\min\left\{y^{\circ}_{t},z^{\circ}_{t},v_{\textup{A}}^{*}\left(x^{\circ}_{t}\right)\right\}\quad\forall t<T.$
    and
    $\min\left\{y^{\circ}_{T+1},z^{\circ}_{T+1},v_{\textup{A}}^{*}\left(x^{\circ}_{T+1}\right)\right\}<\min\left\{y^{\circ}_{T},z^{\circ}_{T},v_{\textup{A}}^{*}\left(x^{\circ}_{T}\right)\right\}.$
    But since $y^{\circ}_{t}$ is non-decreasing and $\min\{z^{\circ}_{t},v_{\textup{A}}^{*}(x^{\circ}_{t})\}$ is non-increasing in $t$, it follows that $\min\{y^{\circ}_{t},z^{\circ}_{t},v_{\textup{A}}^{*}(x^{\circ}_{t})\}$ must achieve its maximal value at the smallest $t$ for which it strictly decreases from $t$ to $t+1$, i.e.
    $\min\left\{\bar{y}_{T},\bar{z}_{T},v_{\textup{A}}^{*}\left(\bar{x}_{T}\right)\right\}$ $=\min\left\{y^{\circ}_{T},z^{\circ}_{T},v_{\textup{A}}^{*}\left(x^{\circ}_{T}\right)\right\}$
    $=\max_{t\in\mathbb{N}}\min\left\{y^{\circ}_{t},z^{\circ}_{t},v_{\textup{A}}^{*}\left(x^{\circ}_{t}\right)\right\}$
    $\geq\max_{t\in\mathbb{N}}\min\left\{r\left(x^{\circ}_{t}\right),z^{\circ}_{t},v_{\textup{A}}^{*}\left(x^{\circ}_{t}\right)\right\}$
    $=\tilde{v}_{\textup{RA}}^{*}(s).$
    where the final equality follows from our choice of $\pi$. Since $\min\{\bar{y}_{t},\bar{z}_{t},v_{\textup{A}}^{*}(\bar{x}_{t})\}$ is non-decreasing in $t$, then
    $\min\{\bar{y}_{t},\bar{z}_{t}\}\geq\min\{\bar{y}_{t},\bar{z}_{t},v_{\textup{A}}^{*}(\bar{x}_{t})\}\geq\min\{\bar{y}_{T},\bar{z}_{T},v_{\textup{A}}^{*}(\bar{x}_{T})\}=\tilde{v}_{\textup{RA}}^{*}(s)\quad\forall t\geq T.$
    Thus
    $v_{\textup{RAA}}^{*}(s)\geq\min\left\{\max_{t\in\mathbb{N}}r\left(\bar{x}_{t}\right),\min_{t\in\mathbb{N}}q\left(\bar{x}_{t}\right)\right\}=\lim_{t\to\infty}\min\{\bar{y}_{t},\bar{z}_{t}\}\geq\tilde{v}_{\textup{RA}}^{*}(s)=v_{\textup{RAA}}^{*}(s),$
    where the final equality follows from Lemma (4). Thus the proof is complete in this case.
  • ($T=\infty$)
    In this case, $\bar{\pi}(\bar{x}_{t},\bar{y}_{t},\bar{z}_{t})=\pi(\bar{x}_{t})$ for all $t\in\mathbb{N}$. Then $\bar{x}_{t}=x^{\circ}_{t}$, $\bar{y}_{t}=y^{\circ}_{t}$, and $\bar{z}_{t}=z^{\circ}_{t}$ for all $t\in\mathbb{N}$. Also $[\bar{x}_{t}]^{+}=x^{\circ}_{t+1}$, $[\bar{y}_{t}]_{\bar{x}_{t}}^{+}=y^{\circ}_{t+1}$, and $[\bar{z}_{t}]_{\bar{x}_{t}}^{+}=z^{\circ}_{t+1}$ for all $t\in\mathbb{N}$. Thus by definition of $T$,
    $\min\left\{y^{\circ}_{t+1},z^{\circ}_{t+1},v_{\textup{A}}^{*}\left(x^{\circ}_{t+1}\right)\right\}\geq\min\left\{y^{\circ}_{t},z^{\circ}_{t},v_{\textup{A}}^{*}\left(x^{\circ}_{t}\right)\right\}\quad\forall t\in\mathbb{N}.$
    Let $T^{\prime}\in\operatorname*{arg\,max}_{t\in\mathbb{N}}\min\left\{y^{\circ}_{t},z^{\circ}_{t},v_{\textup{A}}^{*}\left(x^{\circ}_{t}\right)\right\}$. Then
    $\min\left\{\bar{y}_{T^{\prime}},\bar{z}_{T^{\prime}},v_{\textup{A}}^{*}\left(\bar{x}_{T^{\prime}}\right)\right\}$ $=\min\left\{y^{\circ}_{T^{\prime}},z^{\circ}_{T^{\prime}},v_{\textup{A}}^{*}\left(x^{\circ}_{T^{\prime}}\right)\right\}$
    $=\max_{t\in\mathbb{N}}\min\left\{y^{\circ}_{t},z^{\circ}_{t},v_{\textup{A}}^{*}\left(x^{\circ}_{t}\right)\right\}$
    $\geq\max_{t\in\mathbb{N}}\min\left\{r\left(x^{\circ}_{t}\right),z^{\circ}_{t},v_{\textup{A}}^{*}\left(x^{\circ}_{t}\right)\right\}$
    $=\tilde{v}_{\textup{RA}}^{*}(s).$
    The rest of the proof the follows the same as the previous case with $T$ replaced by $T^{\prime}$.
Corollary 5.
For all $s\in\mathcal{{S}}$, we have $V_{\textup{RAA}}^{*}(s)=v_{\textup{RAA}}^{*}(s)$.
Proof of Theorem 1.
Theorem 1 is now a direct consequence of the previous corollary together with Corollary 4 and Lemma 4. ∎
-C Proof of RR Main Theorem
We first define the value functions, $V_{\textup{R1}}^{*},V_{\textup{R2}}^{*},\tilde{V}_{\textup{R}}^{*},V_{\textup{RR}}^{*}:\mathcal{{S}}\to\mathbb{R}$ by
$V_{\textup{R1}}^{*}(s)$ $=\max_{\pi\in\Pi}\max_{\tau\in\mathbb{N}}r_{1}\left(\xi_{s}^{\pi}(\tau)\right),$
$V_{\textup{R2}}^{*}(s)$ $=\max_{\pi\in\Pi}\max_{\tau\in\mathbb{N}}r_{2}\left(\xi_{s}^{\pi}(\tau)\right),$
$\tilde{V}_{\textup{R}}^{*}(s)$ $=\max_{\pi\in\Pi}\max_{\tau\in\mathbb{N}}r_{\textup{RR}}\left(\xi_{s}^{\mathbf{a}}(\tau)\right),$
$V_{\textup{RR}}^{*}(s)$ $=\max_{\pi\in\Pi}\min\left\{\max_{\tau\in\mathbb{N}}r_{1}\left(\xi_{s}^{\pi}(\tau)\right),\max_{\tau\in\mathbb{N}}r_{2}\left(\xi_{s}^{\pi}(\tau)\right)\right\}.$
We next define the value functions, $v_{\textup{R1}}^{*},v_{\textup{R2}}^{*},\tilde{v}_{\textup{R}}^{*},v_{\textup{RR}}^{*}:\mathcal{{S}}\to\mathbb{R}$, which maximize over action sequences rather than policies:
$v_{\textup{R1}}^{*}(s)$ $=\max_{\mathbf{a}\in\mathbb{{A}}}\max_{\tau\in\mathbb{N}}r_{1}\left(\xi_{s}^{\mathbf{a}}(\tau)\right),$
$v_{\textup{R2}}^{*}(s)$ $=\max_{\mathbf{a}\in\mathbb{{A}}}\max_{\tau\in\mathbb{N}}r_{2}\left(\xi_{s}^{\mathbf{a}}(\tau)\right),$
$\tilde{v}_{\textup{R}}^{*}(s)$ $=\max_{\mathbf{a}\in\mathbb{{A}}}\max_{\tau\in\mathbb{N}}r_{\textup{RR}}\left(\xi_{s}^{\mathbf{a}}(\tau)\right),$
$v_{\textup{RR}}^{*}(s)$ $=\max_{\mathbf{a}\in\mathbb{{A}}}\min\left\{\max_{\tau\in\mathbb{N}}r_{1}\left(\xi_{s}^{\mathbf{a}}(\tau)\right),\max_{\tau\in\mathbb{N}}r_{2}\left(\xi_{s}^{\mathbf{a}}(\tau)\right)\right\},$
where $r_{\textup{RR}}$ is as in Theorem 2. Observe that for each $s\in\mathcal{{S}}$,
$v_{\textup{R1}}^{*}(s)\geq V_{\textup{R1}}^{*}(s),\quad v_{\textup{R2}}^{*}(s)\geq V_{\textup{R2}}^{*}(s),\quad\tilde{v}_{\textup{R}}^{*}(s)\geq\tilde{V}_{\textup{R}}^{*}(s),\quad v_{\textup{RR}}^{*}(s)\geq V_{\textup{RR}}^{*}(s).$
We now prove a series of lemmas that will be useful in the proof of the main theorem.
Lemma 6.
There are $\pi_{1},\pi_{2}\in\Pi$ such that
$v_{\textup{R1}}^{*}(s)=\max_{\tau\in\mathbb{N}}r_{1}\left(\xi_{s}^{\pi_{1}}(\tau)\right)\text{ and }v_{\textup{R2}}^{*}(s)=\max_{\tau\in\mathbb{N}}r_{2}\left(\xi_{s}^{\pi_{2}}(\tau)\right)$
for all $s\in\mathcal{{S}}$.
Proof.
We will just prove the result for $v_{\textup{R1}}^{*}(s)$ since the other result follows identically. For each $s\in\mathcal{{S}}$, let $\tau_{s}$ be the smallest element of $\mathbb{N}$ for which
$\max_{\mathbf{a}\in\mathbb{{A}}}r_{1}\left(\xi_{s}^{\mathbf{a}}(\tau_{s})\right)=v_{\textup{R1}}^{*}(s).$
Moreover, for each $s\in\mathcal{{S}}$, let $\mathbf{a}_{s}$ be such that
$r_{1}\left(\xi_{s}^{\mathbf{a}_{s}}(\tau_{s})\right)=v_{\textup{R1}}^{*}(s).$
Let $\pi_{1}\in\Pi$ be given by $\pi_{1}(s)=\mathbf{a}_{s}(0)$. It suffices to show that
$r_{1}\left(\xi_{s}^{\pi_{1}}(\tau_{s})\right)=v_{\textup{R1}}^{*}(s)$ (29)
for all $s\in\mathcal{{S}}$, for in this case, we have
$v_{\textup{R1}}^{*}(s)\geq\max_{\tau\in\mathbb{N}}r_{1}\left(\xi_{s}^{\pi_{1}}(\tau)\right)\geq r_{1}\left(\xi_{s}^{\pi_{1}}(\tau_{s})\right)=v_{\textup{R1}}^{*}(s)\quad\forall s\in\mathcal{{S}}.$
We show (29) holds for each $s\in\mathcal{{S}}$ by induction on $\tau_{s}$. First, suppose that $s\in\mathcal{{S}}$ is such that $\tau_{s}=0$. Then
$r_{1}\left(\xi_{s}^{\pi_{1}}(\tau_{s})\right)=r_{1}(s)=r_{1}\left(\xi_{s}^{\mathbf{a}_{s}}(\tau_{s})\right)=v_{\textup{R1}}^{*}(s).$
For the induction step, let $n\in\mathbb{N}$ and suppose that
$r_{1}\left(\xi_{s}^{\pi_{1}}(\tau_{s})\right)=v_{\textup{R1}}^{*}(s)\quad\forall s\in\mathcal{{S}}\text{ such that }\tau_{s}\leq n.$
Now fix some $x\in\mathcal{{S}}$ such that $\tau_{x}=n+1$. Notice that
$v_{\textup{R1}}^{*}(x)$ $\geq v_{\textup{R1}}^{*}\left(f\left(x,\pi_{1}(x)\right)\right)$
$\geq\max_{\mathbf{a}\in\mathbb{{A}}}r_{1}\left(\xi_{f\left(x,\pi_{1}(x)\right)}^{\mathbf{a}}(n)\right)$
$\geq r_{1}\left(\xi_{f\left(x,\pi_{1}(x)\right)}^{\mathbf{a}_{x}|_{1}}(n)\right)$
$=r_{1}\left(\xi_{x}^{[\pi_{1}(x),\mathbf{a}_{x}|_{1}]}(n+1)\right)$
$=r_{1}\left(\xi_{x}^{\mathbf{a}_{x}}(\tau_{x})\right)$
$=v_{\textup{R1}}^{*}(x),$
so that $v_{\textup{R1}}^{*}\left(f\left(x,\pi_{1}(x)\right)\right)=v_{\textup{R1}}^{*}(x)$ and $\tau_{f(x,\pi_{1}(x))}\leq n$. It suffices to show
$\tau_{f(x,\pi_{1}(x))}=n,$ (30)
for then, by the induction assumption, we have
$r_{1}\left(\xi_{x}^{\pi_{1}}(\tau_{x})\right)=r_{1}\left(\xi_{f\left(x,\pi_{1}(x)\right)}^{\pi_{1}}(n)\right)=v_{\textup{R1}}^{*}\left(f\left(x,\pi_{1}(x)\right)\right)=v_{\textup{R1}}^{*}(x).$
To show (30), assume instead that
$\tau_{f(x,\pi_{1}(x))}<n.$
But
$v_{\textup{R1}}^{*}(x)$ $\geq\max_{\mathbf{a}\in\mathbb{{A}}}r_{1}\left(\xi_{x}^{\mathbf{a}}\left(\tau_{f\left(x,\pi_{1}(x)\right)}+1\right)\right)$
$\geq r_{1}\left(\xi_{x}^{\left[\pi_{1}(x),\mathbf{a}_{f(x,\pi_{1}(x))}\right]}\left(\tau_{f\left(x,\pi_{1}(x)\right)}+1\right)\right)$
$=r_{1}\left(\xi_{f(x,\pi_{1}(x))}^{\mathbf{a}_{f(x,\pi_{1}(x))}}\left(\tau_{f\left(x,\pi_{1}(x)\right)}\right)\right)$
$=v_{\textup{R1}}^{*}\left(f\left(x,\pi_{1}(x)\right)\right)$
$=v_{\textup{R1}}^{*}(x),$
so that
$v_{\textup{R1}}^{*}(x)=\max_{\mathbf{a}\in\mathbb{{A}}}r_{1}\left(\xi_{x}^{\mathbf{a}}\left(\tau_{f\left(x,\pi_{1}(x)\right)}+1\right)\right)$
and thus
$\tau_{x}\leq\tau_{f\left(x,\pi_{1}(x)\right)}+1<n+1,$
giving our desired contradiction. ∎
Corollary 6.
For all $s\in\mathcal{{S}}$, we have $V_{\textup{R1}}^{*}(s)=v_{\textup{R1}}^{*}(s)$ and $V_{\textup{R2}}^{*}(s)=v_{\textup{R2}}^{*}(s)$.
Lemma 7.
There is a $\pi\in\Pi$ such that
$\tilde{v}_{\textup{R}}^{*}(s)=\max_{\tau\in\mathbb{N}}r_{\textup{RR}}\left(\xi_{s}^{\pi}(\tau)\right).$
for all $s\in\mathcal{{S}}$.
Proof.
This lemma follows by precisely the same proof as the previous lemma, with $r_{1}$, $v_{\textup{R1}}^{*}$, and $\pi_{1}$ replaced with $r_{\textup{RR}}$, $\tilde{v}_{\textup{R}}^{*}$, and $\pi$ respectively. ∎
Corollary 7.
For all $s\in\mathcal{{S}}$, we have $\tilde{V}_{\textup{R}}^{*}(s)=\tilde{v}_{\textup{R}}^{*}(s)$.
Lemma 8.
Let $\zeta_{1}:\mathbb{N}\to\mathbb{R}$ and $\zeta_{2}:\mathbb{N}\to\mathbb{R}$. Then
$\sup_{\tau\in\mathbb{N}}$ $\max\left\{\min\left\{\zeta_{1}(\tau),\sup_{\tau^{\prime}\in\mathbb{N}}\zeta_{2}(\tau+\tau^{\prime})\right\},\min\left\{\sup_{\tau^{\prime}\in\mathbb{N}}\zeta_{1}(\tau+\tau^{\prime}),\zeta_{2}(\tau)\right\}\right\}$
$=\min\left\{\sup_{\tau\in\mathbb{N}}\zeta_{1}(\tau),\sup_{\tau\in\mathbb{N}}\zeta_{2}(\tau)\right\}.$
Proof.
We proceed by showing both inequalities corresponding to the above equality hold.
  • ($\leq$)
    Observe that
    $\sup_{\tau\in\mathbb{N}}$ $\max\left\{\min\left\{\zeta_{1}(\tau),\sup_{\tau^{\prime}\in\mathbb{N}}\zeta_{2}(\tau+\tau^{\prime})\right\},\min\left\{\sup_{\tau^{\prime}\in\mathbb{N}}\zeta_{1}(\tau+\tau^{\prime}),\zeta_{2}(\tau)\right\}\right\}$
    $\leq\max\left\{\min\left\{\sup_{\tau\in\mathbb{N}}\zeta_{1}(\tau),\sup_{\tau\in\mathbb{N}}\sup_{\tau^{\prime}\in\mathbb{N}}\zeta_{2}(\tau+\tau^{\prime})\right\},\min\left\{\sup_{\tau\in\mathbb{N}}\sup_{\tau^{\prime}\in\mathbb{N}}\zeta_{1}(\tau+\tau^{\prime}),\sup_{\tau\in\mathbb{N}}\zeta_{2}(\tau)\right\}\right\}$
    $=\min\left\{\sup_{\tau\in\mathbb{N}}\zeta_{1}(\tau),\sup_{\tau\in\mathbb{N}}\zeta_{2}(\tau)\right\}$
  • ($\geq$)
    Fix $\varepsilon>0$. Choose $\tau_{1},\tau_{2}\in\mathbb{N}$ such that $\zeta_{1}(\tau_{1})\geq\sup_{\tau\in\mathbb{N}}\zeta_{1}(\tau)-\varepsilon$ and $\zeta_{2}(\tau_{2})\geq\sup_{\tau\in\mathbb{N}}\zeta_{2}(\tau)-\varepsilon$. Without loss of generality, we can assume $\tau_{1}\leq\tau_{2}$. Then
    $\sup_{\tau\in\mathbb{N}}$ $\max\left\{\min\left\{\zeta_{1}(\tau),\sup_{\tau^{\prime}\in\mathbb{N}}\zeta_{2}(\tau+\tau^{\prime})\right\},\min\left\{\sup_{\tau^{\prime}\in\mathbb{N}}\zeta_{1}(\tau+\tau^{\prime}),\zeta_{2}(\tau)\right\}\right\}$
    $\geq\sup_{\tau\in\mathbb{N}}\min\left\{\zeta_{1}(\tau),\sup_{\tau^{\prime}\in\mathbb{N}}\zeta_{2}(\tau+\tau^{\prime})\right\}$
    $\geq\min\left\{\zeta_{1}(\tau_{1}),\sup_{\tau^{\prime}\in\mathbb{N}}\zeta_{2}(\tau_{1}+\tau^{\prime})\right\}$
    $\geq\min\left\{\zeta_{1}(\tau_{1}),\zeta_{2}(\tau_{2})\right\}$
    $\geq\min\left\{\sup_{\tau\in\mathbb{N}}\zeta_{1}(\tau)-\varepsilon,\sup_{\tau\in\mathbb{N}}\zeta_{2}(\tau)-\varepsilon\right\}$
    $=\min\left\{\sup_{\tau\in\mathbb{N}}\zeta_{1}(\tau),\sup_{\tau\in\mathbb{N}}\zeta_{2}(\tau)\right\}-\varepsilon.$
    But since $\varepsilon>0$ was arbitrary, the desired inequality follows.
Lemma 9.
For each $s\in\mathcal{{S}}$,
$\tilde{v}_{\textup{R}}^{*}(s)=v_{\textup{RR}}^{*}(s).$
Proof.
For each $s\in\mathcal{{S}}$,
$\tilde{v}_{\textup{R}}^{*}(s)=$ $\max_{\mathbf{a}\in\mathbb{{A}}}\max_{\tau\in\mathbb{N}}r_{\textup{RR}}\left(\xi_{s}^{\mathbf{a}}(\tau)\right)$ (31)
$=$ $\max_{\mathbf{a}\in\mathbb{{A}}}\max_{\tau\in\mathbb{N}}\max\left\{\min\left\{r_{1}\left(\xi_{s}^{\mathbf{a}}(\tau)\right),v_{\textup{R2}}^{*}\left(\xi_{s}^{\mathbf{a}}(\tau)\right)\right\},\min\left\{v_{\textup{R1}}^{*}\left(\xi_{s}^{\mathbf{a}}(\tau)\right),r_{2}\left(\xi_{s}^{\mathbf{a}}(\tau)\right)\right\}\right\}$ (32)
$=$ $\max_{\mathbf{a}\in\mathbb{{A}}}\max_{\tau\in\mathbb{N}}\max\bigg{\{}\min\left\{r_{1}\left(\xi_{s}^{\mathbf{a}}(\tau)\right),\max_{\mathbf{a}^{\prime}\in\mathbb{{A}}}\max_{\tau^{\prime}\in\mathbb{N}}r_{2}\left(\xi_{\xi_{s}^{\mathbf{a}}(\tau)}^{\mathbf{a}^{\prime}}(\tau^{\prime})\right)\right\},$
$\quad\qquad\qquad\qquad\min\left\{\max_{\mathbf{a}^{\prime}\in\mathbb{{A}}}\max_{\tau^{\prime}\in\mathbb{N}}r_{1}\left(\xi_{\xi_{s}^{\mathbf{a}}(\tau)}^{\mathbf{a}^{\prime}}(\tau^{\prime})\right),r_{2}\left(\xi_{s}^{\mathbf{a}}(\tau)\right)\right\}\bigg{\}}$
$=$ $\max_{\mathbf{a}\in\mathbb{{A}}}\max_{\tau\in\mathbb{N}}\max\bigg{\{}\min\left\{r_{1}\left(\xi_{s}^{\mathbf{a}}(\tau)\right),\max_{\mathbf{a}^{\prime}\in\mathbb{{A}}}\max_{\tau^{\prime}\in\mathbb{N}}r_{2}\left(\xi_{s}^{[\mathbf{a},\mathbf{a}^{\prime}]_{\tau}}(\tau+\tau^{\prime})\right)\right\},$
$\quad\qquad\qquad\qquad\min\left\{\max_{\mathbf{a}^{\prime}\in\mathbb{{A}}}\max_{\tau^{\prime}\in\mathbb{N}}r_{1}\left(\xi_{s}^{[\mathbf{a},\mathbf{a}^{\prime}]_{\tau}}(\tau+\tau^{\prime})\right),r_{2}\left(\xi_{s}^{\mathbf{a}}(\tau)\right)\right\}\bigg{\}}$
$=$ $\max_{\mathbf{a}\in\mathbb{{A}}}\max_{\tau\in\mathbb{N}}\max_{\mathbf{a}^{\prime}\in\mathbb{{A}}}\max\bigg{\{}\min\left\{r_{1}\left(\xi_{s}^{\mathbf{a}}(\tau)\right),\max_{\tau^{\prime}\in\mathbb{N}}r_{2}\left(\xi_{s}^{[\mathbf{a},\mathbf{a}^{\prime}]_{\tau}}(\tau+\tau^{\prime})\right)\right\},$
$\quad\qquad\qquad\qquad\qquad\min\left\{\max_{\tau^{\prime}\in\mathbb{N}}r_{1}\left(\xi_{s}^{[\mathbf{a},\mathbf{a}^{\prime}]_{\tau}}(\tau+\tau^{\prime})\right),r_{2}\left(\xi_{s}^{\mathbf{a}}(\tau)\right)\right\}\bigg{\}}$
$=$ $\max_{\mathbf{a}\in\mathbb{{A}}}\max_{\tau\in\mathbb{N}}\max_{\mathbf{a}^{\prime}\in\mathbb{{A}}}\max\bigg{\{}\min\left\{r_{1}\left(\xi_{s}^{[\mathbf{a},\mathbf{a}^{\prime}]_{\tau}}(\tau)\right),\max_{\tau^{\prime}\in\mathbb{N}}r_{2}\left(\xi_{s}^{[\mathbf{a},\mathbf{a}^{\prime}]_{\tau}}(\tau+\tau^{\prime})\right)\right\},$
$\quad\qquad\qquad\qquad\qquad\min\left\{\max_{\tau^{\prime}\in\mathbb{N}}r_{1}\left(\xi_{s}^{[\mathbf{a},\mathbf{a}^{\prime}]_{\tau}}(\tau+\tau^{\prime})\right),r_{2}\left(\xi_{s}^{[\mathbf{a},\mathbf{a}^{\prime}]_{\tau}}(\tau)\right)\right\}\bigg{\}}$ (33)
$=$ $\max_{\mathbf{a}\in\mathbb{{A}}}\max_{\tau\in\mathbb{N}}\max\bigg{\{}\min\left\{r_{1}\left(\xi_{s}^{\mathbf{a}}(\tau)\right),\max_{\tau^{\prime}\in\mathbb{N}}r_{2}\left(\xi_{s}^{\mathbf{a}}(\tau+\tau^{\prime})\right)\right\},$
$\quad\qquad\qquad\qquad\min\left\{\max_{\tau^{\prime}\in\mathbb{N}}r_{1}\left(\xi_{s}^{\mathbf{a}}(\tau+\tau^{\prime})\right),r_{2}\left(\xi_{s}^{\mathbf{a}}(\tau)\right)\right\}\bigg{\}}$ (34)
$=$ $\max_{\mathbf{a}\in\mathbb{{A}}}\min\left\{\max_{\tau\in\mathbb{N}}r_{1}\left(\xi_{s}^{\mathbf{a}}(\tau)\right),\max_{\tau\in\mathbb{N}}r_{2}\left(\xi_{s}^{\mathbf{a}}(\tau)\right)\right\}$ (35)
$=$ $v_{\textup{RR}}^{*}(s),$
where the equality between 31 and 32 follows from Corollary 6, the equality between 33 and 34 follows from Lemma 3, and the equality between 34 and 35 follows from Lemma 8. ∎
Before the next lemma, we need to introduce two last pieces of notation. First, we let $\overline{\Pi}$ be the set of augmented policies $\bar{\pi}:\mathcal{{S}}\times\mathcal{{Y}}\times\mathcal{{Z}}\to\mathcal{{A}}$, as in the previous section, but where
$\mathcal{{Y}}=\left\{r_{1}(s)\mid s\in\mathcal{{S}}\right\}\quad\text{and}\quad\mathcal{{Z}}=\left\{r_{2}(s)\mid s\in\mathcal{{S}}\right\}.$
Next, given $s\in\mathcal{{S}}$, $y\in\mathcal{{Y}}$, $z\in\mathcal{{Z}}$, and $\bar{\pi}\in\overline{\Pi}$, we let $\bar{\xi}_{s}^{\bar{\pi}}:\mathbb{N}\to\mathcal{{S}}$, $\bar{\eta}_{s}^{\bar{\pi}}:\mathbb{N}\to\mathcal{{Y}}$, and $\bar{\zeta}_{s}^{\bar{\pi}}:\mathbb{N}\to\mathcal{{Z}}$, be the solution of the evolution
$\bar{\xi}_{s}^{\bar{\pi}}(t+1)$ $=f\left(\bar{\xi}_{s}^{\bar{\pi}}(t),\bar{\pi}\left(\bar{\xi}_{s}^{\bar{\pi}}(t),\bar{\eta}_{s}^{\bar{\pi}}(t),\bar{\zeta}_{s}^{\bar{\pi}}(t)\right)\right),$
$\bar{\eta}_{s}^{\bar{\pi}}(t+1)$ $=\max\left\{r_{1}\left(\bar{\xi}_{s}^{\bar{\pi}}(t+1)\right),\bar{\eta}_{s}^{\bar{\pi}}(t)\right\},$
$\bar{\zeta}_{s}^{\bar{\pi}}(t+1)$ $=\max\left\{r_{2}\left(\bar{\xi}_{s}^{\bar{\pi}}(t+1)\right),\bar{\zeta}_{s}^{\bar{\pi}}(t)\right\},$
for which $\bar{\xi}_{s}^{\bar{\pi}}(0)=s$, $\bar{\eta}_{s}^{\bar{\pi}}(0)=r_{1}(s)$, and $\bar{\xi}_{s}^{\bar{\pi}}(0)=r_{2}(s)$.
Lemma 10.
There is a $\bar{\pi}\in\overline{\Pi}$ such that
$v_{\textup{RR}}^{*}(s)=\min\left\{\max_{\tau\in\mathbb{N}}r_{1}\left(\bar{\xi}_{s}^{\bar{\pi}}(\tau)\right),\max_{\tau\in\mathbb{N}}r_{2}\left(\bar{\xi}_{s}^{\bar{\pi}}(\tau)\right)\right\}$
for all $s\in\mathcal{{S}}$.
Proof.
By Lemmas 6 and 7 together with Corollary 6, we can choose $\pi,\theta_{1},\theta_{2}\in\Pi$ such that
$v_{\textup{R1}}^{*}(s)$ $=\max_{\tau\in\mathbb{N}}r_{1}\left(\xi_{s}^{\theta_{1}}(\tau)\right)\quad\forall s\in\mathcal{{S}},$
$v_{\textup{R2}}^{*}(s)$ $=\max_{\tau\in\mathbb{N}}r_{2}\left(\xi_{s}^{\theta_{2}}(\tau)\right)\quad\forall s\in\mathcal{{S}},$
$\tilde{v}_{\textup{R}}^{*}(s)$ $=\max_{\tau\in\mathbb{N}}\max\left\{\min\left\{r_{1}\left(\xi_{s}^{\pi}(\tau)\right),v_{\textup{R2}}^{*}\left(\xi_{s}^{\pi}(\tau)\right)\right\},\min\left\{r_{2}\left(\xi_{s}^{\pi}(\tau)\right),v_{\textup{R1}}^{*}\left(\xi_{s}^{\pi}(\tau)\right)\right\}\right\}\quad\forall s\in\mathcal{{S}}.$
Define $\bar{\pi}\in\overline{\Pi}$ by
$\bar{\pi}(s,y,z)=\begin{cases}\pi(s)&\max\{y,z\}<\tilde{v}_{\textup{R}}^{*}(s)\\ \theta_{1}(s)&\max\{y,z\}\geq\tilde{v}_{\textup{R}}^{*}(s)\text{ and }y\leq z,\\ \theta_{2}(s)&\max\{y,z\}\geq\tilde{v}_{\textup{R}}^{*}(s)\text{ and }y>z.\end{cases}$
Now fix some $s\in\mathcal{{S}}$. For all $t\in\mathbb{N}$, set $\bar{x}_{t}=\bar{\xi}_{s}^{\bar{\pi}}(t)$, $\bar{y}_{t}=\bar{\eta}_{s}^{\bar{\pi}}(t)=\max_{\tau\leq t}r_{1}(\bar{x}_{\tau})$, and $\bar{z}_{t}=\bar{\zeta}_{s}^{\bar{\pi}}(t)=\max_{\tau\leq t}r_{2}(\bar{x}_{\tau})$, and also set $x^{\circ}_{t}=\xi_{s}^{\pi}(t)$. It suffices to show
$v_{\textup{RR}}^{*}(s)\leq\min\left\{\max_{\tau\in\mathbb{N}}r_{1}\left(\bar{x}_{\tau}\right),\max_{\tau\in\mathbb{N}}r_{2}\left(\bar{x}_{\tau}\right)\right\},$ (36)
since the reverse inequality is immediate. We proceed in three steps.
  1. 1.
    We claim there exists a $t\in\mathbb{N}$ such that $\max\left\{r_{1}(\bar{x}_{t}),r_{2}(\bar{x}_{t})\right\}\geq\tilde{v}_{\textup{R}}^{*}(\bar{x}_{t})$.
    Suppose otherwise. Then $\bar{\pi}(\bar{x}_{t},\bar{y}_{t},\bar{z}_{t})=\pi(\bar{x}_{t})$ so that $\bar{x}_{t}=x^{\circ}_{t}$ for all $t\in\mathbb{N}$. Thus
    $\max_{t\in\mathbb{N}}\max\left\{r_{1}(\bar{x}_{t}),r_{2}(\bar{x}_{t})\right\}$ $<\max_{t\in\mathbb{N}}\tilde{v}_{\textup{R}}^{*}(\bar{x}_{t})$
    $=\tilde{v}_{\textup{R}}^{*}(s)$
    $=\max_{\tau\in\mathbb{N}}\max\left\{\min\left\{r_{1}\left(x^{\circ}_{\tau}\right),v_{\textup{R2}}^{*}\left(x^{\circ}_{\tau}\right)\right\},\min\left\{r_{2}\left(x^{\circ}_{\tau}\right),v_{\textup{R1}}^{*}\left(x^{\circ}_{\tau}\right)\right\}\right\}$
    $=\max_{\tau\in\mathbb{N}}\max\left\{\min\left\{r_{1}\left(\bar{x}_{\tau}\right),v_{\textup{R2}}^{*}\left(\bar{x}_{\tau}\right)\right\},\min\left\{r_{2}\left(\bar{x}_{\tau}\right),v_{\textup{R1}}^{*}\left(\bar{x}_{\tau}\right)\right\}\right\}$
    $\leq\max_{\tau\in\mathbb{N}}\max\left\{r_{1}(\bar{x}_{\tau}),r_{2}(\bar{x}_{\tau})\right\},$
    providing the desired contradiction.
  2. 2.
    Let $T$ be the smallest element of $\mathbb{N}$ for which
    $\max\left\{r_{1}(\bar{x}_{T}),r_{2}(\bar{x}_{T})\right\}\geq v_{\textup{R}}^{*}(\bar{x}_{T}),$
    which must exist by the previous step, and let $T^{\prime}$ be the smallest element of $\mathbb{N}$ for which
    $\max\left\{\min\left\{r_{1}\left(x^{\circ}_{T^{\prime}}\right),v_{\textup{R2}}^{*}\left(x^{\circ}_{T^{\prime}}\right)\right\},\min\left\{r_{2}\left(x^{\circ}_{T^{\prime}}\right),v_{\textup{R1}}^{*}\left(x^{\circ}_{T^{\prime}}\right)\right\}\right\}=\tilde{v}_{\textup{R}}^{*}(s),$
    which must exist by our choice of $\pi$. We claim $T^{\prime}\geq T$.
    Suppose otherwise. Since $\bar{x}_{t}=x^{\circ}_{t}$ for all $t\leq T$, then in particular $\bar{x}_{T^{\prime}}=x^{\circ}_{T^{\prime}}$, so that
    $\max\left\{\min\left\{r_{1}\left(\bar{x}_{T^{\prime}}\right),v_{\textup{R2}}^{*}\left(\bar{x}_{T^{\prime}}\right)\right\},\min\left\{r_{2}\left(\bar{x}_{T^{\prime}}\right),v_{\textup{R1}}^{*}\left(\bar{x}_{T^{\prime}}\right)\right\}\right\}=\tilde{v}_{\textup{R}}^{*}(s).$
    But then
    $\max\{r_{1}(\bar{x}_{T^{\prime}}),r_{2}(\bar{x}_{T^{\prime}})\}\geq\tilde{v}_{\textup{R}}^{*}(s)\geq\tilde{v}_{\textup{R}}^{*}(\bar{x}_{T^{\prime}}).$
    By our choice of $T$, we then have $T\leq T^{\prime}$, creating a contradiction.
  3. 3.
    It follows from the previous step that
    $\tilde{v}_{\textup{R}}^{*}(\bar{x}_{T})=\tilde{v}_{\textup{R}}^{*}(x^{\circ}_{T})=\tilde{v}_{\textup{R}}^{*}(s).$
    By our choice of $T$, there are two cases: $r_{1}(\bar{x}_{T})\geq\tilde{v}_{\textup{R}}^{*}(\bar{x}_{T})$ and $r_{2}(\bar{x}_{T})\geq\tilde{v}_{\textup{R}}^{*}(\bar{x}_{T})$. We assume the first case and prove the desired result, with case two following identically. To reach a contradiction, assume
    $r_{2}(\bar{x}_{t})<\tilde{v}_{\textup{R}}^{*}(\bar{x}_{T})\quad\forall t\in\mathbb{N}.$
    But then $\bar{\pi}(\bar{x}_{t},\bar{y}_{t},\bar{z}_{t})=\theta_{2}(\bar{x}_{t})$ for all $t\geq T$, so $v_{\textup{R2}}^{*}(\bar{x}_{T})=\max_{t\geq T}r_{2}(\bar{x}_{t})<\tilde{v}_{\textup{R}}^{*}(\bar{x}_{T})\leq\tilde{v}_{\textup{R}}^{*}(s)$. Thus $r_{2}(x^{\circ}_{T^{\prime}})\leq v_{\textup{R2}}^{*}(x^{\circ}_{T^{\prime}})\leq v_{\textup{R2}}^{*}(x^{\circ}_{T})=v_{\textup{R2}}^{*}(\bar{x}_{T})<\tilde{v}_{\textup{R}}^{*}(s)$. It follows that
    $\max\left\{\min\left\{r_{1}\left(x^{\circ}_{T^{\prime}}\right),v_{\textup{R2}}^{*}\left(x^{\circ}_{T^{\prime}}\right)\right\},\min\left\{r_{2}\left(x^{\circ}_{T^{\prime}}\right),v_{\textup{R1}}^{*}\left(x^{\circ}_{T^{\prime}}\right)\right\}\right\}<\tilde{v}_{\textup{R}}^{*}(s),$
    contradicting our choice of $T^{\prime}$.
    Thus $r_{2}(\bar{x}_{t})\geq\tilde{v}_{\textup{R}}^{*}(\bar{x}_{T})=\tilde{v}_{\textup{R}}^{*}(s)$ for some $t\in\mathbb{N}$ and also $r_{1}(\bar{x}_{T})\geq\tilde{v}_{\textup{R}}^{*}(\bar{x}_{T})=\tilde{v}_{\textup{R}}^{*}(s)$, so that (36) must hold by Lemma 9.
Corollary 8.
For all $s\in\mathcal{{S}}$, we have $V_{\textup{RR}}^{*}(s,r_{1}(s),r_{2}(s))=v_{\textup{RR}}^{*}(s)$.
Proof of Theorem 2.
The proof of this theorem immediately follows from the previous corollary together with Corollary 7 and Lemma 9. ∎
-D Proof of Optimality Theorem
Proof of Theorem 3.
The inequalities in both lines of the theorem follow from the fact that for each $\pi\in\Pi$, we can define a corresponding augmented policy $\bar{\pi}\in\overline{\Pi}$ by
$\bar{\pi}(s,y,z)=\pi(s)\quad\forall s\in\mathcal{{S}},y\in\mathcal{{Y}},z\in\mathcal{{Z}},$
in which case $V_{\textup{RAA}}^{\pi}(s)=V_{\textup{RAA}}^{\bar{\pi}}(s)$ and $V_{\textup{RR}}^{\pi}(s)=V_{\textup{RR}}^{\bar{\pi}}(s)$ for each $s\in\mathcal{{S}}$. Note that in general, we cannot define a corresponding policy for each augmented policy, so the reverse inequality does not generally hold (see Figure 2 for intuition regarding this fact).
The equalities in both lines of the theorem are simply restatements of Lemma 5 and Lemma 9. ∎
-E The SRABE and its Policy Gradient
Proof of Proposition 1.
We here closely follow the proof of Theorem 3 in [21], which itself modifies the proofs of the Policy Gradient Theorems in Chapter 13.2 and 13.6 [22]. We only make the minimal modifications required to adapt the PPO algorithm developed previously for the SRBE to on for the SRABE.
$\nabla_{\theta}\tilde{V}_{\textup{RAA}}^{\pi_{\theta}}(s)=$ $\nabla_{\theta}\left(\sum_{a\in\mathcal{{A}}}\pi_{\theta}(a|s)\tilde{Q}_{\textup{RAA}}^{\pi_{\theta}}(s,a)\right)$
$=$ $\sum_{a\in\mathcal{{A}}}\Big{(}\nabla_{\theta}\pi_{\theta}(a|s)\tilde{Q}_{\textup{RAA}}^{\pi_{\theta}}(s,a)$
$~{}~{}\qquad+\pi_{\theta}(a|s)\nabla_{\theta}\min\left\{\max\left\{\tilde{V}_{\textup{RAA}}^{\pi}\left(f(s,a)\right),r_{\textup{RAA}}(s)\right\},q(s)\right\}\Big{)}$
$=$ $\sum_{a\in\mathcal{{A}}}\Big{(}\nabla_{\theta}\pi_{\theta}(a|s)\tilde{Q}_{\textup{RAA}}^{\pi_{\theta}}(s,a)$
$~{}~{}\qquad+\pi_{\theta}(a|s)\left[q(s)<\tilde{V}_{\textup{RAA}}\left(f(s,a)\right)<r_{\textup{RAA}}(s)\right]\nabla_{\theta}\tilde{V}_{\textup{RAA}}^{\pi}\left(f(s,a)\right)\Big{)}$ (37)
$=$ $\sum_{s^{\prime}\in\mathcal{{S}}}\left[\left(\sum_{k=0}^{\infty}\mathrm{Pr}(s\to s^{\prime},k,\pi)\right)\sum_{a\in\mathcal{{A}}}\nabla_{\theta}\pi_{\theta}(a|s^{\prime})\tilde{Q}_{\textup{RAA}}^{\pi_{\theta}}(s^{\prime},a)\right]$ (38)
$=$ $\sum_{s^{\prime}\in\mathcal{{S}}}\left[\left(\sum_{k=0}^{\infty}\mathrm{Pr}(s\to s^{\prime},k,\pi)\right)\sum_{a\in\mathcal{{A}}}\pi_{\theta}(a|s^{\prime})\frac{\nabla_{\theta}\pi_{\theta}(a|s^{\prime})}{\pi_{\theta}(a|s^{\prime})}\tilde{Q}_{\textup{RAA}}^{\pi_{\theta}}(s^{\prime},a)\right]$
$=$ $\sum_{s^{\prime}\in\mathcal{{S}}}\left[\left(\sum_{k=0}^{\infty}\mathrm{Pr}(s\to s^{\prime},k,\pi)\right)\mathbb{E}_{a\sim\pi_{\theta}(s^{\prime})}\left[\nabla_{\theta}\ln\pi_{\theta}(a|s^{\prime})\tilde{Q}_{\textup{RAA}}^{\pi_{\theta}}(s^{\prime},a)\right]\right]$
$\propto$ $~{}\mathbb{E}_{s^{\prime}\sim d_{\pi}^{\prime}(s)}\mathbb{E}_{a\sim\pi_{\theta}(s^{\prime})}\left[\nabla_{\theta}\ln\pi_{\theta}(a|s^{\prime})\tilde{Q}_{\textup{RAA}}^{\pi_{\theta}}(s^{\prime},a)\right],$
where the equality between (37) and (38) comes from rolling out the term $\nabla_{\theta}\tilde{V}_{\textup{RAA}}^{\pi}\left(f(s,a)\right)$ (see Chapter 13.2 in [22] for details), and where $\mathrm{Pr}(s\to s^{\prime},k,\pi)$ is the probability that under the policy $\pi$, the system is in state $s^{\prime}$ at time $k$ given that it is in state $s$ at time 00. ∎
Note, Proposition 1 is vital to updating the actor in Algorithm 1.
-F The DO-HJ-PPO Algorithm
In this section, we outline the details of our Actor-Critic algorithm DO-HJ-PPO beyond the details given in Algorithm 1.
-F1 The special Bellman updates and the corresponding GAEs
Akin to previous HJ-RL policy algorithms, namely RCPO [24], RESPO [20] and RCPPO [21], DO-HJ-PPO fundamentally depends on the discounted HJ Bellman updates [18]. To solve the RAA and RR problems with the special rewards defined in Theorems 1 & 2, DO-HJ-PPO utilizes the Reach, Avoid and Reach-Avoid Bellman updates, given by
$B^{\gamma}_{R}[Q\mid r](s,a)$ $=(1-\gamma)r(s)+\gamma\max\left\{r(s),Q(s,a)\right\},$ (39)
$B^{\gamma}_{A}[Q\mid q](s,a)$ $=(1-\gamma)q(s)+\gamma\min\left\{q(s),Q(s,a)\right\},$ (40)
$B^{\gamma}_{RA}[Q\mid r,q](s,a)$ $=(1-\gamma)\min\left\{r(s),q(s)\right\}+\gamma\min\left\{q(s),\max\left\{r(s),Q(s,a)\right\}\right\}.$ (41)
To improve our algorithm, we incorporate the Generalized Advantage Estimate corresponding to these Bellman equations in the updates of the Actors. As outlined in Section A of [21], the GAE may be defined with a reduction function corresponding to the appropriate Bellman function which will be applied over a trajectory roll-out. We generalize the Reach GAE definition given in [21] to propose a Reach-Avoid GAE (the Avoid GAE is simply the flip of the Reach GAE) as all will be used in DO-HJ-PPO algorithm for either RAA or RR problems. Consider a reduction function $\phi_{RA}^{(n)}:\mathbb{R}^{n}\to\mathbb{R}$, defined by
$\phi_{RA}^{(n)}(x_{1},x_{2},x_{3},\dots,x_{2n+1})$ $=\phi_{RA}^{(1)}(x_{1},x_{2},\phi_{RA}^{(n-1)}(x_{3},\dots,x_{2n+1})),$ (42)
$\phi_{RA}^{(1)}(x,y,z)$ $=(1-\gamma)\min\left\{x,y\right\}+\gamma\min\left\{y,\max\left\{x,z\right\}\right\}.$ (43)
The $k$-step Reach-Avoid Bellman advantage $A_{RA}^{\pi(k)}$ is then given by,
$A_{RA}^{(k)}(s)=\phi_{RA}^{(n)}\left(r(s_{t}),q(s_{t}),\dots,r(s_{t+k-1}),q(s_{t+k-1}),V^{(}s_{t+k})\right)-V^{(}s_{t+k}).$ (44)
We may then define the Reach-Avoid GAE $A_{RA}^{\lambda}$ as the $\lambda$-weighted sum over the advantage functions
$A_{RA}^{\lambda}(s)=\frac{1}{1-\lambda}\sum_{k=1}^{\infty}\lambda^{k}A^{(k)}_{RA}(s)$ (45)
which may be approximated over any finite trajectory sample. See [21] for further details.
-F2 Modifications from standard PPO
To address the RAA and RR problems, DO-HJ-PPO introduces several key modifications to the standard PPO framework [23]:
Additional actor and critic networks are introduced to represent the decomposed objectives.
Rather than learning the decomposed objectives separately from the composed objective, DO-HJ-PPO optimizes all objectives simultaneously. This design choice is motivated by two primary factors: (i) simplicity and minor computational speed-up, and (ii) coupling between the decomposed and composed objectives during learning.
The decomposed trajectories are initialized using states sampled from the composed trajectory, we refer to as coupled resets.
While it is possible to estimate the decomposed objectives independently—i.e., prior to solving the composed task—this approach might lead to inaccurate or irrelevant value estimates in on-policy settings. For example, in the RAA problem, the decomposed objective may prioritize avoiding penalties, while the composed task requires reaching a reward region without incurring penalties. In such a case, a decomposed policy trained in isolation might converge to an optimal strategy within a reward-irrelevant region, misaligned with the overall task. Empirically, we observe that omitting coupled resets causes DO-HJ-PPO to perform no better than standard baselines such as CPPO, whereas their inclusion significantly improves performance.
The special RAA and RR rewards are defined using the decomposed critic values and updated using their corresponding Bellman equations.
This procedure is directly derived from our theoretical results (Theorems 1 and 2), which establish the validity of using modified rewards within the respective RA and R Bellman frameworks. These rewards are used to compute the composed critic target as well as the actor’s GAE. In Algorithm 1, this process is reflected in the critic and actor updates corresponding to the composed objective.
-G DDQN Demonstration
As described in the paper, we demonstrate the novel RAA and RR problems in a 2D $Q$-learning problem where the value function may be observed easily. We juxtapose these solitons with those of the previously studied RA and R problems which consider more simple objectives. To solve all values, we employ the standard Double-Deep $Q$ learning approach (DDQN) [25] with only the special Bellman updates.
-G1 Grid-World Environment
The environment is taken from [19] and consists of two dimensions, $s=(x,y)$, and three actions, $a\in\{\text{left},\text{straight},\text{right}\}$, which allow the agent to maneuver through the space. The deterministic dynamics of the environment are defined by constant upward flow such that,
$f((x_{i},y_{i}),a_{i})=\begin{cases}(x_{i-1},y_{i+1})\quad&a_{i}=\text{left}\\ (x_{i},y_{i+1})\quad&a_{i}=\text{straight}\\ (x_{i+1},y_{i+1})\quad&a_{i}=\text{right}\end{cases}$ (46)
and if the agent reaches the boundary of the space, defined by $x\geq|2|$, $y\leq-2$ and $y\geq 10$, the trajectory is terminated. The 2D space is divided into $80\times 120$ cells which the agent traverses through.
In the RA and RAA experiments, the reward function $r$ is defined as the negative signed-distance function to a box with dimensions $(x_{c},y_{c},w,h)=(0,4.5,2,1.5)$, and thus is negative iff the agent is outside of the box. The penalty function $q$ is defined as the minimum of three (positive) signed distance functions for boxes defined at $(x_{c},y_{c},w,h)=(\pm 0.75,3,1,1)$ and $(x_{c},y_{c},w,h)=(0,6,2.5,1)$, and thus is positive iff the agent is outside of all boxes.
In the R and RR experiments, one or two rewards are used. In the R experiment, the reward function $r$ is defined as the maximum of two negative signed-distance function of boxes with dimensions $(x_{c},y_{c},w,h)=(\pm 1.25,0,0.5,2)$, and thus is negative iff the agent is outside of both boxes. In the RR experiment, the rewards $r_{1}$ and $r_{2}$ are defined as the negative signed distance functions of the same two boxes independently, and thus are positive if the agent is in one box or the other respectively.
-G2 DDQN Details
As per our theoretical results in Theorems 1 and 2, we may now perform DDQN to solve the RAA and RR problems with solely the previously studied Bellman updates for the RA [19] and R problems [18]. We compare these solutions with those corresponding to the RA and R problems without the special RAA and RR targets, and hence solve the previously posed problems. For all experiments, we employ the same adapted algorithm as in [19], with no modification of the hyper-parameters given in Table I.
TABLE I: Hyperparameters for DDQN Grid World
DDQN hyperparameters Values
Network Architecture MLP
Numbers of Hidden Layers 2
Units per Hidden Layer 100, 20
Hidden Layer Activation Function tanh
Optimizer Adam
Discount factor $\gamma$ 0.9999
Learning rate 1e-3
Replay Buffer Size 1e5 transitions
Replay Batch Size 100
Train-Collect Interval 10
Max Updates 4e6
-H Baselines
In both RAA and RR problems, we employ Constrained PPO (CPPO) [2] as the major baseline as it can handle secondary objectives which are reformulated as constraints. The algorithm was not designed to minimize its constraints necessarily but may do so in attempting to satisfy them. As a novel direction in RL, few algorithms have been designed to optimize max/min accumulated costs and thus CPPO serves as the best proxy. Below we also include a naively decomposed STL algorithm to offer some insight into direct approaches to optimizing the max/min accumulated reward.
-H1 CPPO Baselines
Although CPPO formulations do not directly consider dual-objective optimization, the secondary objective in RAA (avoid penalty) or overall objective in RR (reach both rewards) may be transformed into constraints to be satisfied of a surrogate problem. For the RAA problem, this may be defined as
$\operatorname{max}_{\pi}\mathbb{E}_{\pi}\left[\sum_{t}^{\infty}\gamma^{t}\max_{t^{\prime}\leq t}r(s_{t^{\prime}}^{\pi})\right]\quad\text{ s.t. }\quad\min_{t}q(s_{t}^{\pi})\geq 0.$ (47)
For the RR problem, one might propose that the fairest comparison would be to formulate the surrogate problem in the same fashion, with achievement of both costs as a constraint, such that
$\operatorname{max}_{\pi}\mathbb{E}_{\pi}\left[\sum_{t}^{\infty}\gamma^{t}\min\left\{\max_{t^{\prime}\leq t}r_{1}(s_{t^{\prime}}^{\pi}),\max_{t^{\prime}\leq t}r_{2}(s_{t^{\prime}}^{\pi})\right\}\right]\quad\text{ s.t. }\quad\min\left\{\max_{t}r_{1}(s_{t}^{\pi}),\max_{t}r_{2}(s_{t}^{\pi})\right\}\geq 0,$ (48)
which we define as variant 1 (CPPO-v1). Empirically, however, we found this formulation to be the poorest by far, perhaps due to the abundance of the non-smooth combinations. We thus also compare with more naive formulations which relax the outer minimizations to summation in the reward
$\operatorname{max}_{\pi}\mathbb{E}_{\pi}\left[\sum_{t}^{\infty}\gamma^{t}\max_{t^{\prime}\leq t}r_{1}(s_{t^{\prime}}^{\pi})+\max_{t^{\prime}\leq t}r_{2}(s_{t^{\prime}}^{\pi})\right]\quad\text{ s.t. }\quad\min\left\{\max_{t}r_{1}(s_{t}^{\pi}),\max_{t}r_{2}(s_{t}^{\pi})\right\}\geq 0,$ (49)
which we define as variant 2 (CPPOv2), and additionally, in the constraint
$\operatorname{max}_{\pi}\mathbb{E}_{\pi}\left[\sum_{t}^{\infty}\gamma^{t}\max_{t^{\prime}\leq t}r_{1}(s_{t^{\prime}}^{\pi})+\max_{t^{\prime}\leq t}r_{2}(s_{t^{\prime}}^{\pi})\right]\quad\text{ s.t. }\quad\max_{t}r_{1}(s_{t}^{\pi})+\max_{t}r_{2}(s_{t}^{\pi})\geq 0,$ (50)
which we define as variant 3 (CPPOv3). This last approach, although naive and seemingly unfair, vastly outperforms the other variants in the RR problem.
-H2 STL Baselines
In contrast with constrained optimization, one might also incorporate the STL methods, which in the current context simply decompose and optimize the independent objectives. For the RAA problem, the standard RA solution serves as a trivial STL baseline since we may attempt to continuously attempt to reach the solution while avoiding the obstacle. In the RR case, we define a decomposed STL baseline (DSTL) which naively solves both R problems, and selects the one with lower value to achieve first.
-I Details of RAA & RR Experiments: Hopper
The Hopper environment is taken from Gym [26] and [21]. In both RAA and RR problems, we define rewards and penalties based on the position of the Hopper head, which we denote as $(x,y)$ in this section.
In the RAA task, the reward is defined as
$r(x,y)=\sqrt{\|x-2\|+|y-1.4\|}-0.1$ (51)
to incentive the Hopper to reach its head to the position at $(x,y)=(2,1.4)$. The penalty $q$ is defined as the minimum of signed distance functions to a ceiling obstacle at $(1,0)$, wall obstacles at $x>2$ and $x<0$ and a floor obstacle at $y<0.5$. In order to safely arrive at high reward (and always avoid the obstacles), the Hopper thus must pass under the ceiling and not dive or fall over in the achievement of the target, as is the natural behavior.
In the RR task, the first reward is defined again as
$r_{1}(x,y)=\sqrt{\|x-2\|+|y-1.4\|}-0.1$ (52)
to incentive the Hopper to reach its head to the position at $(x,y)=(2,1.4)$, and the second reward as
$r_{2}(x,y)=\sqrt{\|x-0\|+|y-1.4\|}-0.1$ (53)
to incentive the Hopper to reach its head to the position at $(x,y)=(0,1.4)$. In order to achieve both rewards, the Hopper must thus hop both forwards and backwards without crashing or diving.
In all experiments, the Hopper is initialized in the default standing posture at a random $x\in[0,2]$ so as to learn a position-agnostic policy. The DO-HJ-PPO parameters used to train these problems can be found in Table II.
TABLE II: Hyperparameters for Hopper Learning
Hyperparameters for DO-HJ-PPO Values
Network Architecture MLP
Units per Hidden Layer 256
Numbers of Hidden Layers 2
Hidden Layer Activation Function tanh
Entropy coefficient Linear Decay 1e-2 $\rightarrow$ 0
Optimizer Adam
Discount factor $\gamma$ Linear Anneal 0.995 $\rightarrow$ 0.999
GAE lambda parameter 0.95
Clip Ratio 0.2
Actor Learning rate Linear Decay 3e-4 $\rightarrow$ 0
Reward/Cost Critic Learning rate Linear Decay 3e-4 $\rightarrow$ 0
Add’l Hyperparameters for CPPO
$K_{P}$ 1
$K_{I}$ 1e-4
$K_{D}$ 1
-J Details of RAA & RR Experiments: F16
The F16 environment is taken from [21], including a F16 fighter jet with a 26 dimensional observation. The jet is limited to a flight corridor with up to $2000$ relative position north ($x_{PN}$), $1200$ relative altitude ($x_{H}$), and ±$\pm 500$ relative position east ($x_{PE}$).
In the RAA task, the reward is defined as
$r(x,y)=\frac{1}{5}|x_{PN}-1500|-50$ (54)
to incentivize the F16 to fly through the geofence defined by the vertical slice at $1500$ relative position north. The penalty $q$ is defined as the minimum of signed distance functions to geofence (wall) obstacles at $x_{PN}>2000$ and $|x_{PE}|>500$ and a floor obstacle at $x_{H}<0$. In order to safely arrive at high reward (and always avoid the obstacles), the F16 thus must fly through the target geofence and then evade crashing into the wall directly in front of it.
In the RR task, the rewards are defined as
$r_{1}(x_{PN},x_{H})=\frac{1}{5}\sqrt{\|x_{PN}-1250\|+|y-850\|}-30$ (55)
and
$r_{2}(x_{PN},x_{H})=\frac{1}{5}\sqrt{\|x_{PN}-1250\|+|y-350\|}-30$ (56)
to incentive the F16 to reach both low and high-altitude horizontal cylinders. In order to achieve both rewards, the F16 must thus aggressively pitch, roll and yaw between the two targets.
In all experiments, the F16 is initialized with position $x_{PN}\in[250,750],x_{H}\in[300,900],x_{PE}\in[-250,250]$ and velocity in $v\in[200,450]$. Additionally, the roll, pitch, and yaw are initialized with ±$\pm\pi/16$ to simulate a variety of approaches to the flight corridor. Further details can be found in [21]. The DO-HJ-PPO parameters used to train these problems can be found in Table III.
TABLE III: Hyperparameters for F16 Learning
Hyperparameters for DO-HJ-PPO Values
Network Architecture MLP
Units per Hidden Layer 256
Numbers of Hidden Layers 2
Hidden Layer Activation Function tanh
Entropy coefficient Linear Decay 1e-2 $\rightarrow$ 0
Optimizer Adam
Discount factor $\gamma$ Linear Anneal 0.995 $\rightarrow$ 0.999
GAE lambda parameter 0.95
Clip Ratio 0.2
Actor Learning rate Linear Decay 1e-3 $\rightarrow$ 0
Reward/Cost Critic Learning rate Linear Decay 1e-3 $\rightarrow$ 0
Add’l Hyperparameters for CPPO
$K_{P}$ 1
$K_{I}$ 1e-4
$K_{D}$ 1
-K Broader Impacts
This paper touches on advancing fundamental methods for Reinforcement Learning. In particular, this work falls into the class of methods designed for Safe Reinforcement Learning. Methods in this class are primarily intended to prevent undesirable behaviors in virtual or cyber-physical systems, such as preventing crashes involving self-driving vehicles or potentially even unacceptable speech among chatbots. It is an unfortunate truth that safe learning methods can be repurposed for unintended use cases, such as to prevent a malicious agent from being captured, but the authors do not foresee the balance of potential beneficial and malicious applications of this method to be any greater than other typical methods in Safe Reinforcement Learning.
-L Acknowledgments
We would like to thank Kai-Chieh Hsu and Oswin So for sharing their code bases which were fundamental to the success of the project.